A Bibliography of Case Studies in Mathematical Modeling—Ecology, Physiology, and Cell Biology

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.22

Title word cross-reference

+ [393, 172, 190]. ^{+2} [341]. ^{2+} [331, 280]. ^{cdc2} [406, 276]. ^{rum1} [406]. _{3} [380, 284]. α [201]. β [563]. • [216].

-Cells [563]. -gated [380]. -motoneurones [201]. -trisphosphate [380].

17th [544]. 1991 [547].

21st [515]. 24th [515].

3 [296].

5-hydydroxytryptophan [201]. 5-trisphophate-receptor-based [341]. 5-Trisphosphate [331, 310].

5-trisphosphate-gated [317]. 5-trisphosphate-induced [253, 290, 280].

6th [524].

9th [524].

ability [114]. abnormality [212]. abnormally [415]. Absence [410]. abstract [303]. abundance [498]. across [498]. actin [382, 297, 313, 279, 159]. acting [406]. action [301, 241]. activate [306]. activating [79]. activation [379, 486, 276]. active [20]. activity [44, 226, 264]. actomyosin [261, 373]. actomyosin-S1 [373]. acute [201]. Adam [471]. Adaptation [543, 326, 485, 155, 506, 398]. adapting [92]. adaptive [468, 233]. adenosine [216]. ADP [241]. Advanced [75]. Advances [542]. Affairs [400]. afferent [92]. after [201]. Age [570, 95, 421, 472, 324]. Age-Structured [95, 570]. aggregates [166, 199, 389, 432, 492, 282, 283, 319]. aggregation [468]. aging [457]. agonist [328, 341]. agonist-stimulated [341]. airway [401]. algae [511]. algal [150, 353]. Algebra [56]. allele [129]. alleles [88, 388]. alleviating [397]. allocation [324]. alterations [50]. amoeba [1]. among [555, 487, 211, 74]. AMP [238]. analog [318]. Analysis [547, 223, 326, 52, 336, 251, 76, 403, 77, 185, 360, 446, 111, 102, 243, 275, 452, 507, 371]. Analytic [117]. analytical [369, 155]. analyzed [43]. Anaphase [349]. Ancestral [387, 557]. ancestry [422]. anisotropic [53, 196]. annihilation [415, 299]. Annual [544]. anomalous [147]. antibody [331]. aortic [394]. Aplysia [72, 264]. apnea [401]. Appearance [165]. Application [265, 527, 320, 6]. Applications [27, 534, 362, 176, 517, 454]. Applied [433, 207, 248]. appositions [157]. Approach [352, 397]. approaches [537]. Approximation [527, 96, 91]. April [524]. aquatic [305]. aqueous [385]. Arbacia [3]. architecture [243, 394, 14]. area [385]. Arnold [386]. Arrhythmia [192]. arrhythmias [181]. Art [528]. Artificial [337, 70, 367]. aspect [154]. Aspects [158, 121]. assay [261]. assimilation [194, 316]. assisted [131, 151]. associated [350]. Association [400, 313]. asymmetry [210]. Asymptotic [76, 185, 243]. ATP [297, 241, 216, 159]. atrial [152, 283]. attachment [51]. attainability [312]. AUTO [163]. Automatic [401]. autonomous [236]. axon [98, 20].

bacteria [425, 482]. bacterial [506]. bank [421, 513, 246]. based [254, 537, 289, 419, 332, 399, 507, 341, 398]. basis [442]. be [81, 311]. beating [166, 432, 492]. Bedside [532]. beg [191]. beg-hedging [191]. Behavior [130, 117, 262, 348, 183, 492, 266, 41, 153]. Behavioral [186, 352, 59]. behaviour [148]. Bell [284]. Bell-shaped [284]. Bet [553]. bets [244]. between [57, 296, 297, 101, 373]. bi [84]. bi-domain [84]. biased [480, 481]. biconvection [466, 467]. bidomain [348, 489, 490, 371, 455]. Bifurcation [127, 433, 362, 163, 102]. Bifurcations [48, 128, 294, 465, 136, 181, 386, 108, 115, 237, 138]. Bilateral [397]. bilayer [224, 385, 91, 171]. bilayers [94, 197]. bileaflet [151]. binding [373]. Bioconvection [569, 511, 230, 214]. biofluid [365]. Biogeography [28]. Biological [546, 40, 81, 542, 52, 309, 33, 104, 136, 115, 118, 447, 367, 21]. Biology [377, 125, 534, 544, 551, 558, 434, 239, 531, 573, 520, 522, 538, 219, 400, 229, 42]. Biomass [556]. Biomechanics [302]. Biotic [541, 477]. Biphasic [453, 375]. Bistability [201]. bistable [320]. bivalve [189]. Block [331, 274]. Blood [414]. Body [114]. boundary [390]. boundedness [225]. brain [503, 396]. breakdown [224, 165, 197, 91, 171]. Breaks [192]. breeding [502]. bridge [51, 148, 364, 159]. bridges [65]. Bruno [524]. budget [189]. Buffalo [515]. bundle [157]. Bursting [492, 72]. butterfly [151].

C [84]. Ca [380, 331, 280, 341]. Ca-induced [289]. cable [77]. cadmium [82]. Calcium [545, 286, 255, 290, 296, 304, 245, 563, 253, 284, 380, 287, 79, 323, 203, 488, 299, 329, 210, 361, 310, 452, 311, 174, 399, 454, 507, 314, 175]. calcium-gated [284]. calcium-response [284]. Calculations [91]. can [81, 311]. Canada [544]. canine [196, 231, 34]. capacitance [385]. Capacitive [413]. captive [502]. Carbon [189]. Cardiac [567, 101, 335, 340, 192, 532, 322, 257, 127, 181, 108, 115, 348, 486, 301, 170, 501, 274, 275, 453, 157, 399, 74]. cardiomyocyte [461]. Cardiovascular [519]. Cascade [344]. Case [555, 573]. cat [201]. catheter [339]. cations [380]. caused [512]. Causes [292]. causing [213]. cdc}2 [173]. cdc}25 [172]. Cell [377, 134, 558, 434, 561, 531, 525, 573, 560, 532, 219, 379, 94, 468, 78, 166, 199, 54, 389, 391, 432, 492, 42, 393, 357, 63, 111, 112, 36, 173, 157, 456, 510, 461, 282, 283, 319]. Cells [563, 108, 203, 150, 168, 351, 37, 311, 463]. Cellular [520, 453]. cerebellar [67]. cerebellum [284, 380, 317]. cerevisiae} [112]. chain [303]. Change [541, 385, 356]. Changes [8, 179, 351, 512]. changing [403, 242]. channel [507]. channels [284, 380, 399, 317]. Chaos [198, 322, 127, 115, 69, 440, 272, 396]. Chaotic [162, 483, 183]. Chapter [277]. characteristic [212]. Characterization [245, 216]. Checkpoints [456]. Chemical [526, 510]. chemicals [226]. Chemostat [451]. chemotactic [425]. Chemotaxis [43, 205, 506]. chick [389, 432, 492, 319]. chip [413]. Chlamydomonas [466, 467, 229]. chloride [210]. chromosome [410, 427, 458]. chromosomes [422]. Chronic [67, 82, 397]. ciliate [50]. circle [379, 254]. Circuits [264]. circus [257]. cladocera [179, 114, 132]. clamp [367, 368]. clamping [21]. classification [181]. cleavage [126]. Climate [541]. Clinical [350]. Clocks [198, 118]. Clonal [99, 179, 324]. Closed [156, 193, 185, 441, 374]. Closed-loop [156, 441, 374]. Closing [379]. closure [152]. Co [149]. Co-operative [149]. coagonist [290]. coalescence [508]. coalescent [298]. Coalescents [409]. coastal [189]. code [412, 392]. coeruleus [50]. Coexistence [89, 142, 105, 99]. coli} [43, 354]. collection [504]. collective [137, 168]. combined [340]. Combining [332]. common [422].

Communication [53]. communities [537, 300]. community [90]. comparative [395]. comparison [3]. competing [353]. Competition [419, 451, 225, 381, 90]. Competitive [93, 105, 114]. Complex [405, 209, 237, 269, 282, 117]. Compliance [424]. comprehensive [229, 360]. computation [283]. Computational [365, 137]. Computer [65, 131, 151, 368, 206, 152, 447]. Computer-assisted [131, 151]. Computer-generated [368]. Computing [528]. concentration [194, 425, 316, 341]. concentrative [149]. concept [33]. concepts [517, 295]. conditional [59]. conditions [87, 390, 218]. conductance [317]. conductances [367, 368]. conduction [6, 380]. conductor [202, 185]. Conference [530, 544, 524]. conflict [164]. consensual [13]. consequences [7, 147]. conservation [502]. conservatism [484]. considerations [514]. constant [374]. constants [2, 5, 38, 373]. constraints [484]. constriction [270]. constructed [264]. Construction [223]. contaminant [459]. contaminants [499]. continuation [163]. continuum [273]. contracting [382]. Contraction [343, 158, 11, 161, 55, 8, 12, 51, 148]. Control [78, 29, 98, 561, 562, 193, 401, 86, 414, 87, 54, 429, 69, 41, 360, 446, 63, 271, 156, 172, 173, 143, 281]. controlled [361]. Controlling [322, 440, 272, 396]. convection [22]. copepodology [246]. copepods [260]. core [185]. core-conductor [185]. cortical [392]. couplage [11]. coupled [318]. Coupling [220, 261]. Course [62]. covariance [505]. Cracking [412]. created [160]. creating [491]. critical [85, 165, 266]. critique [370]. Croonian [285]. Cross [148, 65, 8, 51, 364, 159]. Cross-bridge [148, 51, 364, 159]. cross-bridges [65]. cross-striations [8]. Crossbridge [565, 241, 216]. crustacea [179, 132]. crustacean [5, 246]. cultivation [189]. cultures [17]. Current [547, 32, 6, 275]. currents [170]. curve [81, 31]. curves [284, 109, 72, 342]. cutaneous [339]. Cycle [525, 560, 379, 134, 386, 54, 109, 42, 393, 357, 111, 112, 173, 456, 510, 159]. Cycles [566]. Cyclic [187, 238]. Cyclin [126, 445, 276, 379]. cycling [212, 270]. cyclins [379]. Cytoplasmic [41, 134, 203, 175]. cytosolic [245, 310].

D [84]. D-C [84]. damped [405]. Daphnia [194, 179, 17, 324, 99, 82, 267, 132, 504, 10, 218, 316, 259, 122]. Data [443, 557, 363, 508]. deaths [213]. decade [167]. December [547]. decerebrate [201]. Deep [503]. defibrillate [501]. defibrillating [170]. defibrillation [160, 195, 202, 288, 486, 206, 430, 101, 32, 275, 340, 375, 462]. defibrillation-strength [462]. deformations [516]. degenerative [457]. Delay [527, 320, 146, 263, 492, 235, 236, 355, 440, 496]. delay-differential [235, 236, 440]. Delayed [562, 405, 474, 475, 183, 209, 269]. delivery [441]. Demographic [484]. Demography [428, 476]. demyelinative [212]. densities [291, 211, 279]. dependence [439, 220]. dependent [70]. Depolarization [301, 288]. depth [230]. Derivation [489]. descent [96]. describing [84]. description [6]. design [131, 151, 441]. destroyed [126]. destruction [117]. destructive [445]. detective [460]. Determination [363]. determines [394]. Deterministic [207]. development [3, 425]. devices [437]. diabetes [70]. diapause [372]. diapausing [421]. diastole [34]. Dictyostelium [155, 468, 238, 398]. did [460]. difference [465, 262]. differences [66, 324]. Differential [251, 49, 124, 407, 534, 347, 116, 56, 184, 571, 572, 308, 176, 465, 163, 97, 146, 262, 263, 4, 235, 236, 355, 440, 496, 102, 248]. differential-delay [146]. differential-difference [465, 262]. differing [364]. Diffusion [518, 96, 91]. dilation [270]. Dimensional [192, 43, 53, 196, 323, 231, 183, 206, 265, 303, 83, 339, 371]. dimensions [287]. dimorphisms [164]. diode [320]. Direct [486]. Directional [66]. disc [131]. discharges [92]. discoideum [468, 238, 155, 398]. Discontinuities [135, 170]. Discontinuous [29, 254]. discrete [468]. disease [325, 479, 426, 457]. Diseases [417]. Disorder [529, 180]. dispersal [483, 300, 330, 242]. dispersion [491]. dissociation [313]. distance [297, 159]. Distemper [555, 487]. distinct [153]. Distribution [231, 147, 265, 32, 449]. distributions [88, 275]. divalent [380]. diversity [472, 477]. division [126, 78, 86, 54, 63, 143]. DNA [178, 50, 36, 133, 508, 557, 457]. do [221, 306]. does [416, 485, 501]. dog [152, 14]. dogs [160, 375]. domain [84]. domains [33]. Dormancy [411]. doubling [108, 115]. Down [192, 493]. drift [68]. driven [115]. drug [441]. duration [288, 423]. durch [1]. during [414, 8, 51, 148, 351, 492, 41, 34, 340, 159, 319, 462, 463]. Dynamic [48, 554, 438, 441, 368, 320, 259, 2, 153, 215, 367]. Dynamical [252, 27, 162, 128, 56, 100, 308, 312]. Dynamics [249, 519, 546, 293, 294, 232, 169, 556, 568, 549, 451, 278, 315, 192, 378, 220, 254, 405, 134, 108, 295, 117, 147, 483, 168, 389, 139, 82, 268, 121, 237, 269, 498, 305, 307, 361, 335, 365, 113, 504, 395, 10, 399, 507]. dysrhythmias [115].

each [126]. ear [437]. early [208]. eastern [325]. Ecological [549, 200, 428]. ecologist [246]. Ecology [552, 402, 548, 417, 186, 352, 573, 179, 345, 537, 259, 497]. ecosystem [366, 395]. ecosystems [537]. edibility [234]. edible [419]. edulis} [226, 459]. Effect [255, 170, 218, 316, 280, 57, 194, 226, 477, 23, 321]. Effective [390]. Effects [378, 430, 152, 24, 324, 234, 240, 241, 455, 218]. efficiencies [316]. efficiency [57]. egg [79, 421, 513]. eggs [126, 134, 421, 3, 331, 306, 174]. elastic [516]. Electric [567, 91, 94, 469, 165, 351, 206, 391, 491, 501, 37]. Electrical [224, 38, 195, 197, 348, 5, 429, 171, 337]. electrically [282]. electrocardiogram [80]. Electrochemical [192]. electrode [455]. electrodes [160, 101, 32, 339]. Electrofusion [531]. electronic [437]. Electrophysiology [532]. electroporating [431, 455]. Electroporation [469, 531, 385, 430]. electrostimulation [397]. element [305, 275]. Elementary [249]. Elements [433]. Elevations [255]. embryonic [166, 389, 283, 319]. embryos [360]. EMBS [544]. Empirical [293]. encodes [173]. encoding [258, 190, 454]. end [479, 445]. endogenous [72]. endoplasmic [284]. endothelial [203]. Energetic [158]. énergetique [11]. Energy [324, 114, 261]. Engineering [534, 544]. Engineers [75]. enrichment [419]. enrichment-nutrient-based [419]. Entrainment [566, 500, 115, 350]. Environment [46, 403, 25, 383, 256, 388, 89, 242, 449]. Environmental [105, 472, 21, 372, 356]. Environments [553, 278, 315, 543, 321, 381, 145, 420, 167, 147, 268, 438, 439, 495, 330]. Epicardial [375, 160]. Epidemic [443, 423]. Epidemics [464, 555, 470, 423, 479]. epilepsy [67, 59, 281]. Epistatic [86]. epithelial [37]. Epizootiology [325]. equation [465, 262, 263]. Equations [251, 49, 124, 407, 534, 347, 116, 56, 184, 571, 527, 572, 308, 176, 222, 85, 163, 97, 146, 4, 235, 236, 355, 440, 496, 102, 248]. equilibrium [476, 394]. Ersatz [1]. Escheria [43]. Escherichia [354]. ESS [145, 312]. estimation [279]. Euglena [493]. Eukaryotic [560]. Evaluation [270, 70].

Eve [400, 370, 460]. events [208]. evoked [203]. Evolution [552, 321, 177, 403, 95, 256, 292, 554, 356, 336, 338, 538, 319, 285, 404, 178, 197, 416, 326, 484, 514, 427, 428, 119, 68, 330, 242, 217, 457]. evolutionarily [449]. Evolutionary [221, 402, 352, 354, 300, 133, 247, 345, 106, 484, 244, 312, 460]. Exact [88, 303]. examination [485]. Exchangeability [119, 524]. excitability [329]. Excitable [536, 250, 346, 500, 307, 369, 454]. excitation [30, 196, 6, 274, 453, 506, 461]. excitatory [130]. exclusion [93]. exhibit [264]. exogenous [70]. expansion [180]. Experiment [432]. Experimental [542, 350, 90, 374]. experimentelle [1]. Experiments [40, 135]. explained [311]. explorations [345]. exponents [180]. external [391]. extinction [403, 416, 479, 356]. Extracellular [196, 265]. extracts [360].

factor [134, 349, 208, 510]. familiar [485]. Fathers [427]. favors [483]. Feedback [40, 562, 193, 405, 296, 182, 183, 209, 269, 441]. feeder [132]. feeding [226, 189]. fertilization [174]. fertilized [331]. FES [193]. Fiber [243, 34, 394, 83]. Fiber-architecture [243]. fibers [26, 424, 216, 373]. fibre [5]. fibres [297, 9]. Fibroblasts [296]. field [94, 160, 196, 351, 391, 501, 453]. Fields [128, 567, 202, 469, 206, 339, 340]. filament [55, 159]. filaments [297, 424]. filling [152]. filter [132]. final [470, 521]. Finetti [524]. Finite [275, 230, 303, 133]. firing [98]. First [62, 355]. first-order [355]. Fisher [96]. fission [78, 86, 446, 172, 173]. fitness [404, 132]. Fixed [386, 492, 448]. fjord [366]. fjords [395]. flat [37]. Fluctuating [553, 260, 383, 473, 438, 439, 495, 505]. fluctuations [472]. Fluid [249, 539, 546, 121, 568, 168, 335]. fluids [447]. focusing [150]. following [50]. food [194, 324, 267, 218, 316]. force [408, 39, 61, 444]. forced [250, 136, 138, 307, 454]. forces [384]. forcing [386]. form [17, 185]. formalism [55]. Formation [542, 58, 197, 482, 310]. fortgesetzte [1]. fotpflanzung [1]. foundations [303]. four [287, 395]. fractal [394]. free [79, 175]. Frequency [454, 258, 439]. Freshwater [509, 260, 246]. frog [77, 41, 216]. Frontiers [548, 497]. function [235]. Functional [347, 132, 429]. functions [172]. fundamental [326]. Fura [245]. Fura-2-loaded [245]. fusion [36].

G [398]. G-protein-based [398]. G1 [379, 393]. G2 [379]. gain [235]. galeata [82, 218, 316]. games [312]. gated [284, 380, 317]. Gaussian [204]. Gene [298, 86, 327, 393, 190, 173]. Genealogical [228, 409, 418]. Genealogical-tree [228]. genealogies [298, 327]. Genealogy [120, 141]. general [447]. generalist [132]. generality [332]. generated [368, 339, 340]. generating [65, 327]. Generation [296, 115, 39]. generations [383, 473, 477, 505]. Genetic [54, 167, 63, 403, 383, 472, 473, 68, 442, 505, 370, 217]. Genetical [15, 64]. Genetics [221, 35, 530, 107, 227, 523, 333, 111, 387]. Genostasis [285]. Geometrical [124]. Geometry [157, 104]. Germination [164, 145]. Global [136, 541, 138, 386]. globus [426]. goes [302]. graded [137]. gradient [160, 339, 340]. Gradients [545, 425, 101]. grazing [353]. gross [50]. Group [521]. Growth [230, 267, 122, 17, 259, 242, 214, 218, 459]. guide [229]. guinea [512]. gyrotactic [230, 214, 273].

habitat [321, 147]. hamster [331]. haploid [141]. harbor [487]. harbour [113]. harmonic [405]. head [444]. heads [382]. hearing [437]. Heart [568, 195, 66, 288, 166, 199, 389, 432, 492, 80, 394, 38, 282, 283, 319]. heartbeat [415]. hearts [14, 462]. heat [2, 23]. heavy [279]. Hedging [244, 191]. held [547, 546, 515]. help [106]. Hematopoietic [561]. Hemispheric [210]. hepatocyte [286]. Hepatocytes [245, 310, 314, 175]. herbivore [305]. herbivorous [494]. herbivory [419]. herds [470]. heterogeneous [321, 167, 326]. High [183, 94]. Hilbert [97]. histamine [203]. Histories [336, 338, 495]. History [345, 553, 7, 504]. Hodgkin [85]. hog [14]. homeostasis [70]. homogeneous [32, 103]. Homogenization [359]. homolog [190]. honour [524]. hopf [102]. Hormone [311, 175]. Hormone-induced [311]. hormone-stimulated [175]. Human [529, 400, 178, 410, 422, 203, 427, 209, 236, 16, 457, 458]. hump [31]. Huxley [85]. hybrid [320, 318]. Hydrid [193]. Hydrodynamic [144, 150, 334]. hydrolysed [297]. hydrolysis [373, 159]. hydrophobic [226]. hydydroxytryptophan [201]. hypotheses [498]. hypothesis [370].

Identification [213]. identified [264]. IEEE [544]. II [265, 490, 558, 155, 91]. III [564]. Illinois [547]. immigration [485]. implantable [437]. Implications [326]. Implicit [439]. importance [200]. impulse [66, 73]. Impulses [19]. inactivation [349]. including [102]. incompressible [222]. incorporating [193]. increasing [61]. independent [310]. Individual [537, 168, 332, 413, 240, 499, 504]. Individual-based [537, 332]. individuals [504]. Individuen [1]. induce [453]. Induced [245, 253, 469, 289, 290, 258, 430, 311, 280, 159, 342]. inducer [172]. inedible [419]. infarction [350]. infection [408]. Infectious [417, 479]. Inference [557, 387]. Inferring [508]. infinite [88]. infinitely [228, 418]. infinitely-many-sites [228, 418]. influence [485, 92]. Information [515, 440, 33]. inhibition [110, 139, 92]. inhibitor [406]. inhibitory [72]. initiated [379, 349]. initiating [369]. Initiation [47]. inlet [189]. inner [437]. Inositol [253, 380, 255, 331, 317, 290, 210, 310, 174, 280, 341]. input [24, 72]. Ins [284]. Insight [235, 117]. Insights [448]. InsP [380]. Instabilities [346, 87]. insulin [70]. insulin-dependent [70]. intact [360]. intensive [189]. interacting [447]. Interaction [567, 301]. Interactions [541, 86, 491, 234]. Interannual [200]. intercellular [231]. interconnecting [318]. Interference [9, 57]. internal [206]. International [530, 544, 524, 527]. interpenetrating [33]. Interpretation [223, 8]. intracardiac [339]. Intracellular [255, 329, 208, 371]. intragenic [129]. intraluminal [380]. intravenous [201]. Introduction [252, 539, 35, 162, 107, 227, 347, 186, 333, 572, 538, 176, 251, 357]. Invariants [345]. Invasibility [225]. investigate [364]. Investigations [293]. Ionic [389]. ip-controlled [361]. ip-sensitive [507]. IRIS [215]. Irregular [212, 108, 182]. irregularity [110]. ischemic [84]. Island [28]. Isochrons [60]. isolated [461]. isometric [26].

Jedi [435]. July [546]. junction [413]. junctional [37]. junctions [231, 371].

Kin [106, 439]. kinase [406, 190, 173]. Kinetic [313, 363, 510]. Kinetics [373, 364, 399]. know [221]. kurz [10].

laboratory [229]. laevis} [299, 134, 329]. Lakes [344, 291, 211]. landscapes [326]. Large [516, 120, 320, 469, 119]. laser [320]. later [167]. latipes [79]. layer [230]. leaflets [394]. learning [429]. Lecture [285]. Leeds [546]. left [243, 34, 83]. length [26]. level [240]. levels [464, 498, 459]. Life [345, 553, 198, 336, 338, 7, 267, 495, 504]. Life-History [553, 504]. lifetime [91]. light [209, 235, 236, 266, 496, 157, 13]. like [137]. Limit [566, 386, 109]. limitation [324]. Limitations [490]. limits [285]. line [20]. Linear [56, 4, 303]. liner [482]. Lines [96]. linked [123]. linking [240]. lipid [94, 224, 197, 91]. little [221]. live [460]. liver [311]. living [9]. load [23]. loaded [245]. local [485]. localized [253]. location [392]. locked [199]. locking [108, 115, 118, 283]. Locomotion [140, 429]. locus [410, 71]. loop [193, 130, 441, 156, 374]. loops [474, 475]. loss [437]. lottery [105]. low [267, 37, 279]. LV [152]. Lyapounov [180].

M [134, 360, 271, 216, 510]. M-phase [271, 510, 360]. M-phase-specific [134]. Machine [429]. Macroscopic [431, 170]. magna} [194, 179, 324, 132]. maintained [473, 505]. maintaining [383, 374]. maintenance [472, 477, 300]. major [479, 423]. mammal [521]. Mammalian [36, 66, 83, 38]. Mammals [547, 521]. Man [45, 374]. Management [540, 502]. Mannheim [527]. many [228, 418]. mapping [375]. maps [254]. Marine [547, 521, 509]. maternal [126]. Mathematical [75, 539, 30, 561, 239, 573, 565, 52, 520, 560, 115, 121, 154, 37]. Mathematics [207, 248]. Matrices [18, 180]. Matrix [223]. maturation [349, 208, 41]. maturation-promoting [349, 208]. may [51]. measurement [194]. Measurements [466, 351, 270, 462]. Mechanical [394, 363]. Mechanics [539, 565, 384, 517]. mechanism [224, 161, 469, 39, 271]. Mechanisms [195, 540, 452, 538, 560, 389, 329]. Mechanochemical [261]. medaka [79]. Media [536, 346, 500, 103]. Medical [535]. Medicine [544]. medium [369]. meeting [546]. meiotic [41]. Membrane [563, 19, 98, 6, 431, 91, 455]. membranes [94, 224, 385, 91, 171, 37]. memory [320, 376]. mendotae} [82]. meromyosin [279]. Messenger [559, 205]. Metapopulation [293, 295]. method [222, 447]. Methods [124, 75, 97, 73]. Mg [280]. micro [58, 230, 480, 481, 149, 214, 273, 334]. micro-organisms [58, 230, 480, 481, 149, 214, 273, 334]. Microbial [525, 451]. microbiology [523]. microscopic [170]. microscopy [9, 157]. Middle [437]. migration [321, 233]. Minimal [258]. Minimum [478]. Mitochondrial [178, 457]. mitosis [379, 406, 358, 445, 36, 190, 47]. mitotic [406, 208, 446, 172]. mitral [121, 131, 151, 152]. Mixed [182, 420, 209, 269]. mixing [464]. mobilization [253]. Model [53, 40, 202, 382, 468, 289, 323, 258, 88, 228, 418, 115, 259, 98, 348, 55, 480, 481, 129, 205, 206, 185, 431, 489, 490, 353, 301, 121, 360, 446, 155, 361, 122, 241, 242, 273, 274, 366, 32, 37, 453, 311, 506, 398, 399, 371, 455, 84, 341, 376]. modeler [456]. Modeling [567, 496, 573, 565]. Modelling [487, 236, 309, 408]. Models [570, 223, 555, 561, 443, 499, 520, 509, 560, 225, 44, 537, 19, 30, 291, 96, 419, 295, 204, 211, 494, 303, 332, 240, 507, 247, 217, 64]. modes [320, 263]. modified [82]. Molecular [219, 377, 292, 329, 434, 520, 522, 538, 442, 446, 400]. molecule [384, 297]. Molecules [560]. molluscs [189]. monitoring [461]. monophasic [453, 375]. monotonic [225]. Montreal [544]. mortality [267]. motile [150]. motoneurones [201]. Motor [302, 450]. Movement [444, 65, 257, 261, 156, 279]. mRNA [126]. multiple [264, 300, 317]. multisite [461]. Multispecies [381]. Multistability [474, 562, 405, 475, 355]. Multitude [262]. Muscle [343, 12, 567, 565, 158, 66, 161, 382, 26, 297, 424, 2, 23, 55, 8, 9, 39, 51, 148, 61, 156, 501, 216, 157, 38, 373, 374, 512]. Muscles [140]. musculaire [11]. muscular [14]. mussels [226, 459]. Mutant [123]. mutants [78]. Mutual [57, 106]. myocardial [84, 463]. myocardium [77, 196, 231, 265]. myocytes [399]. Myosin [450, 382, 384, 297, 444, 363, 216, 313, 279, 159]. myth [400]. Mytilus [226, 459].

nanometre [384]. Natural

[15, 417, 68, 207, 179, 416, 442]. Navier [222]. Negative [190, 405]. neocortical [436]. nerve [19, 30, 5, 6, 20, 73, 281]. nervous [515]. nested [141]. net [218]. network [318, 376]. networks [137, 269, 337]. Neural [137, 269]. neuroelectric [44]. neurology [237]. neuron [413, 130, 318]. Neuronal [254, 412]. neurons [137, 264, 436, 392, 24, 72, 367, 368]. neurooscillator [98]. neuropathy [212]. Neurosurgery [550]. neutral [88, 96, 388, 129]. neutrality [409]. next [379]. niche [484]. Niches [554, 326, 428, 142]. Noise [475, 266, 269, 16]. Non [204, 103, 3, 4]. Non-Gaussian [204]. Non-homogeneous [103]. non-linear [4]. non-nucleate [3]. Nonlinear [518, 128, 536, 232, 184, 251, 465, 262, 263, 138, 389, 236, 355]. normal [80]. North [459, 325]. note [51]. nuclear [41, 143]. nucleate [3]. nucleocytoplasmic [50]. nucleoside [373]. nucleotide [123]. nucleotides [364]. Null [85]. number [64]. Numerical [360, 528, 155, 369, 73]. nutrient [419, 498]. nutrients [268].

observations [295]. observed [323].

obstructive [401]. obtusa} [10]. October [515, 527]. offspring [164]. olivary [318]. One [71, 371, 244, 311, 247, 159]. One- [71]. One-dimensional [371]. one-parameter [247]. onset [235, 266, 271]. oocyte [360]. oocytes [253, 255, 287, 134, 323, 299, 329, 210, 41]. operations [414]. operative [149]. optic [212]. Optical [288, 351, 461, 462]. Optimal [411, 29, 420, 495, 147]. optimization [438]. Optimizing [25]. order [465, 222, 262, 263, 4, 355]. orders [406]. Ordinary [251, 49, 124, 116, 184, 571, 163, 102]. organic [226]. organisms [58, 230, 480, 481, 149, 214, 273, 334]. organization [386, 314]. organs [92]. orientation [34, 83]. origin [200]. origins [400]. orthosis [193]. Oryzias [79]. Oscillating [255]. Oscillation [320, 465, 69, 235, 266, 331]. Oscillations [545, 526, 128, 296, 562, 361, 245, 563, 289, 258, 182, 328, 209, 236, 238, 154, 452, 311, 16, 454, 341]. Oscillator [286, 405, 127, 136, 386, 115, 138]. Oscillators [52, 115, 81, 109, 72, 318]. Oscillatory [263, 529, 310, 262, 337]. outbreak [479]. outbreaks [423]. outcome [234]. Overdrive [432, 74, 282]. Overlapping [477, 383, 473, 505]. oxytactic [482].

P [284]. p25 [406]. p34 [406, 276]. Pacemaker [24, 110]. pacemakers [74]. pain [503, 374]. paired [342]. Paleopopulation [538]. pallidum [397]. pallidus [426]. pancreas [70]. Pancreatic [563]. panoramic [392]. papillary [512]. paradigm [182, 485, 187]. paradox [99]. parameter [247]. parasites [57]. parent [164]. parent-offspring [164]. Parkinson [426]. parkinsonian [397]. Part [552, 55, 265, 489, 490, 558, 564, 174, 155]. Partial [407, 97, 572, 176, 248, 437]. patch [147, 339]. Pathological [87]. pathway [78, 448]. patients [212]. Pattern [58, 482, 542, 466, 467]. Patterns [473, 323, 230, 264, 214]. PDV [487, 378].

Penetrative [22]. perfused [462]. period [288, 108, 115]. Periodic [386, 76, 185, 265, 371, 319]. periodically [127, 136, 108, 115, 199, 138, 500]. persistent [264]. persists [379]. Perspective [404, 554, 456]. Perspectives [232]. Phase [257, 566, 108, 115, 166, 199, 118, 40, 406, 134, 135, 81, 109, 393, 360, 102, 283, 342, 72, 271, 510]. phase-resetting [135]. Phase-response [72]. Phase-Shift [40]. phaseless [60]. phenomena [7, 149, 334, 461]. Phenomenology [224]. phenotype [449]. Phenotypic [177, 513, 204, 68]. Phenylephrine [245]. Phocine [555, 487]. phosphorylation [258, 173]. Photic [277]. photoreceptors [488]. phototactic [511]. Physarum [47, 143]. Physics [534]. physiological [19, 87, 259, 69, 121]. Physiologically [169, 303]. Physiology [535, 573, 564]. Pi [241]. PicoNewton [384]. pig [512]. pivoting [131]. pivoting-disc [131]. place [51]. Plaines [547]. planar [385]. Plane [48, 102]. Plankton [498, 291, 211, 187]. planktonic [99, 153]. plant [300, 305, 142, 246]. plant-herbivore [305]. plants [419]. played [174]. points [386]. polycephalum} [47, 143]. polygenic [217]. polymorphism [410, 473, 167]. pool [311, 341]. poor [268]. Population [223, 35, 411, 548, 147, 554, 82, 333, 305, 113, 549, 10, 278, 315, 509, 217, 378, 179, 7, 17, 387, 260, 478, 483, 494, 497, 240, 499, 242, 504, 141, 133]. Populations [95, 46, 417, 120, 169, 556, 309, 537, 408, 472, 476, 428, 119, 353, 187, 495, 303, 442]. pore [385]. pore-pore [385]. pores [197]. position [151]. positive [296]. Potential [265, 563, 160, 469, 351, 430, 491, 101, 301, 339, 340, 462, 463, 512]. potentially [173]. potentials [84]. Practices [547]. Predation [90, 200]. Predator [268, 220, 353]. Predator-prey [268, 220, 353]. predators [57, 234].

Predicted [40]. Prediction [17]. Predictions [385]. premature [491]. Presence [317]. pressure [401, 414]. Presynaptic [110]. prevent [416]. prey [220, 353, 268, 31]. probabilities [228, 418]. probability [479, 524]. problem [408]. Problems [207, 163, 517]. Proceedings [546, 544, 524, 515, 547]. process [298]. Processes [536, 62]. processing [515, 33]. produced [444]. production [117, 218, 316]. productivity [234]. Products [18, 180]. Professor [524]. Profile [450]. programmes [502]. progression [393]. projection [222]. prolong [288]. promotes [105]. promoting [349, 208, 510]. propagating [346]. propagation [30, 299, 500, 73, 369, 174]. Properties [129, 380, 77, 146, 137, 373, 289]. Proposed [39]. prosthetic [131]. Protein [450, 126, 258, 190, 173, 398]. Proteins [450, 302]. proteolysis [379, 349]. proteus [1]. protozoen [1]. protozoen-Individuen [1]. pulex} [17, 99, 267, 504]. pulse [346, 20]. pulsed [72]. pumped [320]. punctulata [3]. Pupil [21, 235, 236, 266, 212, 270, 496, 13, 16]. pupillary [215]. Purification [208].

Qualitative [571, 102]. Quantitative [221, 226, 530, 107, 227, 446, 505, 403, 6, 204, 442, 171, 452]. quantitative-genetic [403]. quantum [302].

rabbit [288, 424, 216, 462]. Random [125, 18, 180, 480, 481, 68, 449]. randomly [25, 145]. range [94]. rapid [415, 210]. rapidly [382]. Rarity [514, 551]. rate [226, 373]. rather [349]. ratio [50, 220]. Ratio-dependence [220]. re [485]. re-examination [485]. reacting [154]. real [368]. reanalysis [370]. receptor [331, 92, 341]. Recipes [528]. recombination [129, 64]. recording [215]. recordings [288]. recticulum [284].

Recurrent [92, 474, 475, 139, 130]. recycling [305]. reentrant [415, 500]. Reentry [448, 257, 274]. Reflections [555]. Reflex [59, 209, 235, 236, 266, 496, 215, 13]. Reflexes [140, 59]. refractory [288]. regenerationen [1]. regular [182]. Regularly [24]. regulated [173]. regulating [271]. Regulation [488, 393, 280, 380, 36, 190, 510]. regulator [208]. Rehabilitation [547, 502]. Relating [504, 499]. Relation [443, 189]. relationship [74, 373]. relationships [226, 133]. Relaxation [94]. relay [155]. Release [547, 382, 289, 290, 502, 280]. Reliablity [436]. relief [503]. remarks [291]. Repetitive [203, 73, 175, 98]. Replication [561]. repolarization [301, 453]. report [521]. Representations [556, 303]. reproduces [323]. reproduction [25, 259, 267, 218]. required [196]. Rescue [547]. reserves [114]. Resetting [415, 566, 350, 257, 135, 166]. resistive [37]. resonance [250]. Resource [234]. respiration [194]. Response [391, 284, 260, 81, 109, 354, 356, 499, 72, 170, 132, 342]. Responses [463, 137, 210]. resulting [87, 282]. results [155, 73]. Return [435]. reveal [78]. Reversible [197, 165, 171]. revisited [244]. Rhythmogenesis [318]. Rhythms [198, 40, 199, 282, 319]. rich [268]. ring [346, 500]. rises [175]. rod [488]. Role [383, 491]. roles [328]. Rome [524]. rules [324]. rum}1 [393]. ryanodine [399]. ryanodine-sensitive [399].

S [406]. S-phase [406]. S1 [373]. Saccharomyces [112, 522]. Sampling [388, 88]. sarcomere [26]. saturation [469]. Sciences [207]. Scientific [528]. Scientists [75]. Scope [459]. Sea [521, 459, 126, 174, 325, 521, 113]. seal [325, 213, 113]. Seals [555, 378, 487, 502]. searching [57]. Seasonal [46, 200]. Second [530, 559, 465, 222, 78, 262, 263, 205, 4]. second-messenger [205]. second-order [222, 263]. section [521]. sediment [421]. seed [164, 246]. segregating [64]. Selection [15, 71, 473, 106, 416, 260, 68, 439, 505, 247]. senescence [484]. sensitive [399, 507]. sensitivity [488]. Sensory [533, 193]. separate [218]. separation [385]. September [544]. Sequence [557, 458, 508]. sequences [133]. sequential [340]. series [373]. servoanalytical [13]. servomchanism [21]. servomechanism [16]. sets [60]. shaped [284]. Shift [40]. shock [463]. shocks [288, 430, 491, 453, 340, 375, 462, 512]. short [376]. short-term [376]. shortening [2, 23, 61]. show [288]. shown [153]. Signal [289, 559, 258, 155]. Signal-induced [289, 258]. signalling [468, 287, 314]. significance [233]. Silicon [413]. Silicon-neuron [413]. similar [142]. Simple [211, 556, 146, 269, 311]. Simplification [507]. Simulating [348, 20]. Simulation [80, 188, 65, 447, 366]. simulations [448]. Simultaneous [194]. since [471]. Single [384, 295, 245, 391, 130, 444, 153, 340, 341, 342]. single-pool [341]. Single-species [295]. Singularities [40]. sink [428]. sites [228, 418, 64, 123]. size [408, 470, 78, 114, 478, 63, 363, 279]. skeletal [424, 156, 216]. skinned [297, 424]. sleep [401]. Sliding [297, 159, 55, 279]. slow [279]. slowly [92]. slowly-adapting [92]. Smooth [100]. Society [544]. Software [163]. Solution [355, 185, 373]. Some [291, 147, 514, 345]. sons [427]. sound [392]. source [428]. source-sink [428]. Sourcebook [229]. Space [309, 549, 85, 97]. spaced [24]. Spatial [314, 491]. spatially [330]. specialization [321]. Specie [554]. species [114, 295, 142]. specific [134]. specified [126]. speeds [61]. sperm [306]. spike [436]. spikes [203]. spiking [304, 376]. spinal [201]. Spiral [299]. spontaneously [166, 432, 492]. spread [66, 487]. squid [98]. Stability [27, 144, 16, 106, 300, 247]. stable [187, 449]. stage [472]. stage-structured [472]. stages [51]. State [515, 137, 109, 91]. states [19, 317]. Statistical [517]. statistics [524]. steady [109, 91]. steady-state [91]. Stentor [50]. step [363, 279]. steps [384]. stiffness [61]. stimulated [127, 108, 199, 328, 500, 310, 175, 341, 282]. stimulating [170]. Stimulation [426, 215, 67, 413, 351, 429, 265, 492, 503, 281, 319]. stimuli [342]. Stochastic [44, 146, 423, 479, 62, 188, 543, 225, 420]. Stokes [222]. stopped [81]. storage [477]. store [440]. story [460]. strategic [366]. Strategies [411, 553, 145, 420]. straus [194]. strength [288, 462]. stress [354, 82]. stretch [8, 92]. striated [55, 39]. striations [8]. Strong [567, 106]. structural [8]. Structure [443, 250, 179, 226, 409, 231, 12, 90]. structure-activity [226]. Structured [95, 411, 494, 169, 240, 556, 509, 570, 472, 476, 495, 303]. Structured-Population [509]. structures [76]. Studies [573, 94, 261, 350, 152, 80, 36, 313]. study [202, 81, 90, 274, 395, 13, 369]. subfragment [313]. subsamples [141]. substitutions [123]. suggesting [51]. supplement [37]. suppression [432, 74, 282]. surgical [414]. survivorship [421]. suspension [466, 467, 58, 230, 425, 214, 189, 511]. suspensions [482, 238, 273, 334]. swimming [58, 480, 481, 149, 168, 334]. symmetry [345]. symposium [515, 527]. symptoms [397]. synapses [337]. synaptic [24, 72]. synchronous [47]. syncytial [390, 359]. syncytium [53]. synthesis [50, 36]. System [519, 320, 250, 205, 515, 153, 305, 318, 374]. Systems [48, 252, 27, 162, 526, 128, 56, 100, 71, 559, 308, 529, 509, 105, 87, 183, 429, 69, 441, 102, 154, 307, 447, 21, 454]. systole [152, 34].

tachycardia [350]. Tachycardias [540]. take [51]. technique [215]. Techniques [362]. tells [493]. Temporal [529, 154, 314, 179]. temporally [330]. tension [26]. term [376]. Terminal [45]. termination [500]. Terrestrial [509]. test [85]. testability [332]. Tests [498]. Textbook [535]. Their [443, 52, 258, 8, 301]. theorem [102]. Theoretical [402, 293, 55, 333, 542, 283, 19, 30, 514, 448]. theories [484, 12]. Theory [48, 124, 27, 35, 15, 385, 433, 571, 28, 362, 336, 451, 388, 115, 482, 432, 4, 102, 171, 103, 510]. therapeutic [397]. therapy [59]. thermal [354]. thin [424]. thorax [202]. those [137]. Three [206, 153, 83, 339, 192, 43, 196, 114, 231, 265]. Three-dimensional [206, 83, 339, 43, 196, 231, 265]. Time [309, 549, 104, 192, 471, 183]. time-delayed [183]. times [508]. timing [78, 436, 152]. tissue [257, 348, 486, 170]. tissues [390, 359]. tongues [386]. tool [523, 240, 502, 510]. Topological [81]. Topology [566, 342]. torso [32]. trabecular [66, 38]. tracking [43]. traits [204]. trajectories [480, 481]. Transduction [533, 559, 261, 92]. transfer [235]. Transformation [4]. Transient [109, 175]. Transmembrane [512, 469, 351, 430, 491, 462, 463]. transmission [20]. Traveling [526]. traverses [79]. treatment [401]. tree [228, 418]. trickery [303]. Triggered [555]. triphosphate [216]. triphosphates [373]. trisphophate [341]. Trisphosphate [255, 331, 253, 380, 290, 210, 310, 174, 280, 317]. Trophic [344, 234]. Turning [358]. Two [323, 328, 71, 464, 53, 147, 137, 51]. Two-dimensional [323, 53]. Two-locus [71]. two-patch [147]. two-state [137]. types [465, 300, 153]. ubiquitination [445]. Ueber [1]. UK [546]. uncertainty [372]. understanding [510]. Unidirectional [274, 369]. uniform [214, 453]. unifying [484]. Universal [181, 271]. University [546, 515, 527]. Unrooted [418]. unsterblichkeit [1]. until [379]. upon [92]. urchin [126, 174]. use [229, 102, 364]. useful [502]. using [348, 185, 235, 236, 33].

Vagus [281]. valve [121, 152, 394]. Valves [568, 131, 151]. variability [105]. Variable [278, 315, 256, 89, 372]. Variance [177, 505]. Variation [61, 383, 26, 200, 513, 211, 442, 217, 457, 458]. varying [381, 25, 408, 145, 388, 330]. Vasopressin [245]. Vector [128]. ventricle [34]. ventricles [14]. ventricular [77, 166, 350, 351, 301, 243, 83, 340, 375]. ventroposteriolateral [397]. vertebrate [26, 488]. vertical [233]. viable [478]. vibration [103]. Views [547]. Virtual [455]. Virus [487, 213, 555]. vitro [261]. voltage [165]. volume [202]. Vulnerability [369].

Wadden [521, 113]. walk [480, 481]. Walks [125]. wall [243, 83]. Wave [536, 79, 299, 331, 500, 174]. Wavelengths [467, 466]. Waves [545, 526, 192, 323, 361, 310, 452]. weak [180, 247]. wee}1 [190]. within [470]. without [64]. Working [521]. workshop [547]. Wright [96].

X. [329]. Xenopus

[134, 360, 323, 253, 255, 287, 210].

Yeast [522, 379, 78, 86, 54, 446, 63, 172, 173]. York [515].

ZFY [410]. Zone [277]. zooplankton [513, 353, 233, 90, 494].

References

Hartmann:1928:UEU

[1] M. Hartmann. Ueber experimentelle unsterblichkeit von protozoen-individuen. Ersatz der fotpflanzung von amoeba proteus durch fortgesetzte regenerationen. Zool. Jahrb., 45:973–987, 1928.

Hill:1938:HSD

[2] A. V. Hill. The heat of shortening and dynamic constants of muscle. Proceedings of the Royal Society of London. Series B. Biological sciences, 126:136–195, 1938. CODEN PRLBA4. ISSN 0962-8452 (print), 1471-2954 (electronic).

Harvey:1940:CDN

[3] E. B. Harvey. A comparison of the development of nucleate and non-nucleate eggs of Arbacia punctulata. Biological Bulletin, 79:166–187, 1940. CODEN BIBUBX. ISSN 0006-3185.

Levinson:1944:TTN

[4] N. Levinson. Transformation theory of non-linear differential equations of the second order. Annals of Mathematics, 45:723–737, 1944. CODEN AN-MAAH. ISSN 0003-486X (print), 1939-8980 (electronic).

Hodgkin:1946:ECC

[5] A. L. Hodgkin and W. A. H. Rushton. The electrical constants of a crustacean nerve fibre. Proceedings of the Royal Society of London. Series B. Biological sciences, 133:444–479, 1946. CODEN PRLBA4. ISSN 0962-8452 (print), 1471-2954 (electronic).

Hodgkin:1952:QDM

[6] A. L. Hodgkin and A. F. Huxley. A quantitative description of membrane current and its application to conduction and excitation in nerve. Journal of Physiology, 117:500–544, 1952. CODEN JPHYA7. ISSN 0022-3751.

Cole:1954:PCL

[7] L. C. Cole. The population consequences of life history phenomena. Quarterly Review of Biology, 29:103–137, 1954. CO-DEN QRBIAK. ISSN 0033-5770 (print), 1539-7718 (electronic).

Huxley:1954:CCM

[8] H. E. Huxley and J. Hanson. Changes in the cross-striations of muscle during contraction and stretch and their structural interpretation. Nature, 173:973–976, 1954. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Huxley:1954:IML

[9] A. F. Huxley and R. Niedergerke. Interference microscopy of living muscle fibres. Nature, 173:971–973, 1954. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Slobodkin:1954:PDD

[10] L. B. Slobodkin. Population dynamics in Daphnia obtusa kurz. Ecological Monographs, 24:69–88, 1954. CODEN EC-MOAQ. ISSN 0012-9615.

Aubert:1956:CCM

[11] Xavier Aubert. Le couplage énergetique de la contraction musculaire. Éditions Arscia, Brussels, Belgium, 1956. 315 pp. Thèse d’agrégation — Université Catholique de Louvain.

Huxley:1957:MST

[12] A. F. Huxley. Muscle structure and theories of contraction. Progress in Biophysics and Biophysical Chemistry, 7:255–318, 1957. CODEN PBBCAF. ISSN 0096-4174.

Stark:1957:SSC

[13] L. Stark and P. M. Sherman. A servoanalytical study of the consensual pupil reflex to light. Journal of Neurophysiology, 20:17–26, 1957. CODEN JONEA4. ISSN 0022-3077.

Thomas:1957:MAV

[14] C. E. Thomas. The muscular architecture of the ventricles of hog and dog hearts. American Journal of Anatomy, 101:17–57, 1957. CODEN AJANA2. ISSN 0002-9106.

Fisher:1958:GTN

[15] R. A. Fisher. The Genetical Theory of Natural Selection. Dover Publications, Inc., New York, NY, 1958. 291 pp. LCCN QH366 .F5 1958.

Stark:1959:SON

[16] L. Stark. Stability, oscillations, and noise in the human pupil servomechanism. Proceedings of the Institute of Radio Engineers, 47:1925–1939, 1959. CO-DEN PIREAE. ISSN 0096-8390.

Frank:1960:PPG

[17] P. W. Frank. Prediction of population growth form in Daphnia pulex cultures. American Naturalist, 94:357–372, 1960. CODEN AMNTA4. ISSN 0003-0147.

Furstenberg:1960:PRM

[18] H. Furstenberg and H. Kesten. Products of random matrices. Annals of Mathematical Statistics, 31(2):457–469, June 1960. CODEN AASTAD. ISSN 0003-

4851. URL http://projecteuclid.
org/euclid.aoms/1177705909.

FitzHugh:1961:IPS

[19] R. FitzHugh. Impulses and physiological states in theoretical models of nerve membrane. Biophysical Journal, 1:445– 466, 1961. CODEN BIOJAU. ISSN 0006-3495.

Nagumo:1962:APT

[20] J. S. Nagumo, S. Arimoto, and S. Yoshizawa. An active pulse transmission line simulating nerve axon. Proceedings of the Institute of Radio Engineers, 50:2061–2071, 1962. CODEN PIREAE. ISSN 0096-8390.

Stark:1962:ECB

[21] L. Stark. Environmental clamping of biological systems: Pupil servomchanism. Journal of the Optical Society of America. A, Optics, image science, and vision, 52:925–930, 1962. CODEN JOAOD6. ISSN 1084-7529, 0740-3232.

Veronis:1963:PC

[22] G. Veronis. Penetrative convection. Astrophysical Journal, 137:641–663, 1963. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Hill:1964:ELH

[23] A. V. Hill. The effect of load on the heat of shortening muscle. Proceedings of the Royal Society of London. Series B. Biological sciences, 159:297–318, 1964. CO-DEN PRLBA4. ISSN 0962-8452 (print), 1471-2954 (electronic).

Perkel:1964:PNE

[24] D. H. Perkel, J. H. Schulman, T. H. Bullock, G. P. Moore, and J. P. Segundo.

Pacemaker neurons: Effects of regularly spaced synaptic input. Science, 145:61– 63, 1964. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Cohen:1966:ORR

[25] D. Cohen. Optimizing reproduction in a randomly varying environment. Journal of Theoretical Biology, 12:119–129, 1966. CODEN JTBIAP. ISSN 0022-5193.

Gordon:1966:VIT

[26] A. M. Gordon, A. F. Huxley, and F. J. Julian. The variation in isometric tension with sarcomere length in vertebrate muscle fibers. Journal of Physiology, 184:170–192, 1966. CODEN JPHYA7. ISSN 0022-3751.

Bhatia:1967:DSS

[27] N. P. (Nam Parshad) Bhatia and G. P. Szegő. Dynamical Systems: Stability Theory and Applications, volume 35 of Lecture notes in mathematics. Springer-Verlag, Berlin; New York, 1967. 6 + 416 pp. LCCN QA3 .L35 v.35.

MacArthur:1967:TIB

[28] R. H. MacArthur and E. O. Wilson. The Theory of Island Biogeography, volume 1 of Monographs in Population Biology. Princeton University Press, Princeton, NJ, USA, 1967. xi + 203 pp. LCCN QH85 .M3.

Flugge-Lotz:1968:DOC

[29] Irmgard Flügge-Lotz. Discontinuous and Optimal Control. McGraw-Hill, New York, NY, USA, 1968. x + 286 pp. LCCN A402.3 .F53.

FitzHugh:1969:BE

[30] R. FitzHugh. Mathematical models of excitation and propagation in theoretical models of nerve. In Herman P. Schwan et al., editors, Biological Engineering, volume 9 of Inter-university electronic series, pages 1–85. McGraw-Hill, New York, NY, USA, 1969. LCCN R856 .B47.

Rosenzweig:1969:WPC

[31] M. Rosenzweig. Why the prey curve has a hump. American Naturalist, 103: 81–87, 1969. CODEN AMNTA4. ISSN 0003-0147.

Rush:1969:CDD

[32] S. Rush, E. Lepeschkin, and A. Gregoritsch. Current distribution from defibrillation electrodes in a homogeneous torso model. Journal of electrocardiology, 2:331–341, 1969. CODEN JECAB4. ISSN 0022-0736.

Schmitt:1969:BIP

[33] O. H. Schmitt. Biological information processing using the concept of interpenetrating domains. In Leibovic [515], pages 325–331.

Streeter:1969:FOC

[34] D. D. Streeter, Jr., H. M. Spotnitz, D. P. Patel, J. Ross, Jr., and E. H. Sonnenblick. Fiber orientation in the canine left ventricle during diastole and systole. Circulation Research, 24:339–347, 1969. CODEN CIRUAL. ISSN 0009-7330.

Crow:1970:IPG

[35] J. F. Crow and M. Kimura. Introduction to Population Genetics Theory. Burgess Publishing Company, Minneapolis, MN, 1970. ISBN 0-06-356128-X (Harper and Row paperback, 1972 reprint), 0-8087-2910-2. xiv + 591 pp. LCCN QH431 .C76.

Rao:1970:MCF

[36] P. N. Rao and R. T. Johnson. Mammalian cell fusion: studies on the regulation of DNA synthesis and mitosis. Nature, 225:159–164, 1970. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Shiba:1970:EMF

[37] H. Shiba. An electric model for flat epithelial cells with low resistive junctional membranes. A mathematical supplement. Japanese Journal of Applied Physics, 9:1405–1409, 1970. CODEN JJAPA5. ISSN 0021-4922.

Weidmann:1970:ECT

[38] S. Weidmann. Electrical constants of trabecular muscle from mammalian heart. Journal of Physiology, 210:1041– 1054, 1970. CODEN JPHYA7. ISSN 0022-3751.

Huxley:1971:PMF

[39] A. F. Huxley and R. M. Simmons. Proposed mechanism of force generation in striated muscle. Nature, 233:533–538, 1971. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Johnsson:1971:FEB

[40] A. Johnsson and H. G. Karlsson. Biological rhythms: Singularities in phase-shift experiments as predicted from a feedback model. In E. Broda, A. Locker, and H. Springer-Lederer, editors, First European Biophysics Congress, 14–17 Sept. 1971, Baden near Vienna, Austria, pages 263–267. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. LCCN QT34 .E89 1971. Six volumes.

Masui:1971:CCN

[41] Y. Masui and C. L. Markert. Cytoplasmic control of nuclear behavior during meiotic maturation of frog oocytes. Journal of experimental zoology, 177: 129–146, 1971. CODEN JEZOAO. ISSN 0022-104X.

Mitchison:1971:BCC

[42] J. M. Mitchison. The biology of the cell cycle. Cambridge University Press, Cambridge, UK, 1971. ISBN 0-521-08251-X, 0-521-09671-5 (paperback). v + 313 pp. LCCN QH605 .M682 1971.

Berg:1972:CEC

[43] H. C. Berg and D. A. Brown. Chemotaxis in Escheria coli analyzed by threedimensional tracking. Nature, 239: 500–504, 1972. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Cowan:1972:SMN

[44] J. D. Cowan. Stochastic models of neuroelectric activity. In Rice et al. [517], pages 181–182. ISBN 0-226-71115-3. LCCN QC175 .I5 1971. IUPAP Conference on Statistical Mechanics (6th: 1971: Chicago, IL).

Crichton:1972:TM

[45] Michael Crichton. The Terminal Man. Ballantine Books, New York, NY, USA, 1972. ISBN 0-394-44768-9. xi + 247 pp. LCCN PS3553.R48 T4.

Fretwell:1972:PSE

[46] S. D. Fretwell. Populations in a Seasonal Environment. Princeton University Press, Princeton, NJ, USA, 1972.

ISBN 0-691-08105-0, 0-691-08106-9 (paperback). xxiii + 217 pp. LCCN QL752 .F73.

Sachsenmaier:1972:ISM

[47] W. Sachsenmaier, U. Remy, and R. Plattner-Schobel. Initiation of synchronous mitosis in Physarum polycephalum. Experimental Cell Research, 73:41–48, 1972. CODEN ECREAL. ISSN 0014-4827.

Andronov:1973:TBD

[48] A. A. Andronov, E. A. Leontovich, I. I. Gordon, and A. G. Maier. Theory of Bifurcations of Dynamic Systems on a Plane. Wiley, New York, 1973. ISBN 0-470-03194-8. xiv + 482 pp. LCCN QA845 T4 1967.

Arnold:1973:ODE

[49] V. I. (Vladimir Igorevich) Arnold́. Ordinary Differential Equations. MIT Press, Cambridge, MA, 1973. ISBN 0-262-01037-2. ix + 280 pp. LCCN QA372.A713. Translated and edited by Richard A. Silverman.

Frazier:1973:DSF

[50] E. A. J. Frazier. DNA synthesis following gross alterations of the nucleocytoplasmic ratio in the ciliate Stentor coeruleus. Developmental Biology, 34: 77–92, 1973. CODEN DEBIAO. ISSN 0012-1606.

Huxley:1973:NSC

[51] A. F. Huxley. A note suggesting that cross-bridge attachment during muscle contraction may take place in two stages. Proceedings of the Royal Society of London. Series B. Biological sciences, 183:83–86, 1973. CODEN PRLBA4.

ISSN 0962-8452 (print), 1471-2954 (electronic).

Pavlidis:1973:BOT

[52] T. Pavlidis. Biological Oscillators: Their Mathematical Analysis. Academic Press, New York, NY, 1973. ISBN 0-12-547350-8. xiii + 207 pp. LCCN QH527 .P38.

Bukauskas:1974:MTA

[53] F. F. Bukauskas, N. I. Kukushkin , and M. Y. Sakson. Model of a two-dimensional anisotropic syncytium. communication I. Biofizika, 19:712–716, 1974. CODEN BIOFAI. ISSN 0006-3029.

Hartwell:1974:GCC

[54] L. H. Hartwell, J. Culotti, J. R. Pringle, and B. J. Reid. Genetic control of the cell division cycle in yeast. Science, 183:46–51, 1974. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Hill:1974:TFS

[55] T. L. Hill. Theoretical formalism for the sliding filament model of contraction of striated muscle, part I. Progress in Biophysics and Molecular Biology, 28:267– 340, 1974. CODEN PBIMAC. ISSN 0079-6107.

Hirsch:1974:DED

[56] M. W. Hirsch and S. Smale. Differential Equations, Dynamical Systems and Linear Algebra, volume 60 of Pure and applied mathematics. Academic Press, 1974. ISBN 0-12-349550-4. xi + 358 pp. LCCN QA3 .P8 v.60.

Beddington:1975:MIB

[57] J. R. Beddington. Mutual interference between parasites or predators and its effect on searching efficiency. Journal of Animal Ecology, 44:331–340, 1975. CO-DEN JAECAP. ISSN 0021-8790.

Childress:1975:PFS

[58] S. Childress, M. Levandowsky, and E. A. Spiegel. Pattern formation in a suspension of swimming micro-organisms. Journal of Fluid Mechanics, 69:595–613, 1975. CODEN JFLSA7. ISSN 0022-1120.

Forster:1975:REB

[59] Francis M. Forster. Reflex epilepsy, behavioral therapy, and conditional reflexes. Thomas, Springfield, IL, USA, 1975. ISBN 0-398-03614-4. ix + 318 pp. LCCN RC372 .F67.

Guckenheimer:1975:IPS

[60] J. Guckenheimer. Isochrons and phaseless sets. Journal of Mathematical Biology, 1:259–273, 1975. CODEN JM-BLAJ. ISSN 0303-6812.

Julian:1975:VMS

[61] F. J. Julian and M. R. Sollins. Variation of muscle stiffness with force at increasing speeds of shortening. Journal of General Physiology, 66:287–302, 1975. CODEN JGPLAD. ISSN 0022-1295.

Karlin:1975:FCS

[62] S. Karlin and H. M. Taylor. A First Course in Stochastic Processes. Academic Press, New York, NY, second edition, 1975. ISBN 0-12-398552-8. xvi + 557 pp. LCCN QA274 .K371 1975.

Nurse:1975:GCC

[63] P. Nurse. Genetic control of cell size at division in yeast. Nature, 256:547–551, 1975. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Watterson:1975:NSS

[64] G. A. Watterson. On the number of segregating sites in genetical models without recombination. Theoretical Population Biology, 7:256–276, 1975. CODEN TLPBAQ. ISSN 0040-5809.

Brokaw:1976:CSM

[65] C. J. Brokaw. Computer simulation of movement generating crossbridges. Biophysical Journal, 16:1013– 1027, 1976. CODEN BIOJAU. ISSN 0006-3495.

Clerc:1976:DDI

[66] L. Clerc. Directional differences of impulse spread in trabecular muscle from mammalian heart. Journal of Physiology, 255:335–346, 1976. CODEN JPHYA7. ISSN 0022-3751.

Cooper:1976:CCS

[67] I. S. Cooper, I. Amin, M. Riklan, J. M. Waltz, and T. P. Poon. Chronic cerebellar stimulation in epilepsy. Archives of Neurology, 33:559–570, 1976. CODEN ARNEAS. ISSN 0003-9942.

Lande:1976:NSR

[68] R. Lande. Natural selection and random genetic drift in phenotypic evolution. Evolution, 30:314–334, 1976. CO-DEN EVOLAO. ISSN 0014-3820.

Mackey:1977:OCP

[69] M. C. Mackey and L. Glass. Oscillation and chaos in physiological control systems. Science, 197:287–289, 1977. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Mirouze:1977:EEI

[70] J. Mirouze, J. L. Selam, T. C. Pham, and D. Cavadore. Evaluation of exogenous insulin homeostasis by the artificial pancreas in insulin-dependent diabetes. Diabetologia, 13:273–278, 1977. CODEN DBTGAJ. ISSN 0012-186X.

Nagylaki:1977:SOT

[71] T. Nagylaki. Selection in One- and Two-locus Systems, volume 15 of Lecture notes in biomathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. ISBN 0-387-08247-6. vii + 208 pp. LCCN QH455 .N331.

Pinsker:1977:ABN

[72] H. M. Pinsker. Aplysia bursting neurons as endogenous oscillators: I: Phaseresponse curves for pulsed inhibitory synaptic input. Journal of Neurophysiology, 40:527–543, 1977. CODEN JONEA4. ISSN 0022-3077.

Rinzel:1977:RNI

[73] J. Rinzel. Repetitive nerve impulse propagation: Numerical results and methods. In Fitzgibbon, III and Walker [518], pages 186–212. ISBN 0-273-01066-2. LCCN QA274.75 .N66.

Vassalle:1977:RAC

[74] M. Vassalle. The relationship among cardiac pacemakers: Overdrive suppression. Circulation Research, 41:269–277, 1977. CODEN CIRUAL. ISSN 0009-7330.

Bender:1978:AMM

[75] C. M. Bender and S. A. Orszag. Advanced Mathematical Methods for Scientists and Engineers. McGraw-Hill, New York, NY, USA, 1978. ISBN 0-07-004452-X. xiv + 593 pp. LCCN QA371 .B43.

Bensoussan:1978:AAP

[76] Alain Bensoussan, Jacques-Louis Lions, and George Papanicolaou. Asymptotic analysis for periodic structures. North-Holland Publishing Co., Amsterdam, The Netherlands, 1978. ISBN 0-444-85172-0. xxiv + 700 pp. LCCN QA379 .B45.

Chapman:1978:ACP

[77] R. A. Chapman and C. H. Fry. An analysis of cable properties of frog ventricular myocardium. Journal of Physiology, 283:263–282, 1978. CODEN JPHYA7. ISSN 0022-3751.

Fantes:1978:CTC

[78] P. A. Fantes and P. Nurse. Control of the timing of cell division in fission yeast: cell size mutants reveal a second control pathway. Experimental Cell Research, 115:317–329, 1978. CODEN ECREAL. ISSN 0014-4827.

Gilkey:1978:FCW

[79] J. C. Gilkey, L. F. Jaffe, E. B. Ridgeway, and G. T. Reynolds. A free calcium wave traverses the activating egg of the medaka, Oryzias latipes. Journal of Cell Biology, 76:448–466, 1978. CO-DEN JCLBA3. ISSN 0021-9525.

Miller:1978:SSE

[80] W. T. Miller, III and D. B. Geselowitz. Simulation studies of the electrocardiogram. I. the normal heart. Circulation Research, 43:301–315, 1978. CODEN CIRUAL. ISSN 0009-7330.

Kawato:1978:BOC

[81] M. Kawato and R. Suzuki. Biological oscillators can be stopped. topological study of a phase response curve. Biological Cybernetics, 30:241–284, 1978. CO-DEN BICYAF. ISSN 0340-1200.

Marshall:1978:PDD

[82] J. S. Marshall. Population dynamics of Daphnia galeata mendotae as modified by chronic cadmium stress. Journal of the Fisheries Research Board of Canada, 35:461–469, 1978. CODEN JFRBAK. ISSN 0015-296X.

Streeter:1978:TFO

[83] D. D. Streeter, Jr., W. E. Powers, M. A. Ross, and F. Torrent-Guasp. Three-dimensional fiber orientation in the mammalian left ventricular wall. In Baan et al. [519], pages 73–84. ISBN 0-262-18078-2. LCCN RC670.C37.

Tung:1978:BMD

[84] L. Tung. A bi-domain model for describing ischemic myocardial D-C potentials. PhD thesis, MIT, Cambridge, MA, 1978.

Best:1979:NSH

[85] E. N. Best. Null space in the Hodgkin– Huxley equations: A critical test. Biophysical Journal, 27:87–104, 1979. CO-DEN BIOJAU. ISSN 0006-3495.

Fantes:1979:EGI

[86] P. A. Fantes. Epistatic gene interactions in the control of division in fission yeast. Nature, 279:428–430, 1979. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Glass:1979:PCR

[87] L. Glass and M. C. Mackey. Pathological conditions resulting from instabilities in physiological control systems. Annals of the New York Academy of Sciences, 316:214–235, 1979. CODEN ANYAA9. ISSN 0077-8923 (print), 1749-6632 (electronic).

Griffiths:1979:ESD

[88] R. C. Griffiths. Exact sampling distributions from the infinite neutral alleles model. Advances in Applied Probability, 11:326–354, 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

Levins:1979:CVE

[89] R. Levins. Coexistence in a variable environment. American Naturalist, 114: 765–783, 1979. CODEN AMNTA4. ISSN 0003-0147.

Lynch:1979:PCZ

[90] M. Lynch. Predation, competition, and zooplankton community structure: An experimental study. Limnology and Oceanography, 24:253–272, 1979. CO-DEN LIOCAH. ISSN 0024-3590.

Pastushenko:1979:EBB

[91] V. F. Pastushenko, Y. A. Chizmadzhev, and V. B. Arakelyan. Electric breakdown of bilayer lipid membranes: II. calculations of the membrane lifetime in the steady-state diffusion approximation. Bioelectrochemistry and Bioenergetics, 6:53–62, 1979. CODEN BE-BEBP. ISSN 0302-4598.

Vibert:1979:RII

[92] J. F. Vibert, M. Davis, and J. P. Segundo. Recurrent inhibition: Its influence upon transduction and afferent discharges in slowly-adapting stretch receptor organs. Biological Cybernetics, 33: 167–178, 1979. CODEN BICYAF. ISSN 0340-1200.

Armstrong:1980:CE

[93] R. A. Armstrong and R. McGehee. Competitive exclusion. American Naturalist, 115:151–170, 1980. CODEN AM-NTA4. ISSN 0003-0147.

Benz:1980:RSC

[94] R. Benz and U. Zimmermann. Relaxation studies on cell membranes and lipid bilayers in the high electric field range. Bioelectrochemistry and Bioenergetics, 7:723–739, 1980. CODEN BE-BEBP. ISSN 0302-4598.

Charlesworth:1980:EAP

[95] Brian Charlesworth. Evolution in Age-Structured Populations. Cambridge University Press, Cambridge, UK, 1980. ISBN 0-521-29786-9 (paperback), 0-521-23045-4. xiii + 300 pp. LCCN QH455 .C47.

Griffiths:1980:LDD

[96] R. C. Griffiths. Lines of descent in the diffusion approximation of neutral Wright–Fisher models. Theoretical Population Biology, 17:37–50, 1980. CO-DEN TLPBAQ. ISSN 0040-5809.

Gustafson:1980:PDE

[97] Karl E. Gustafson. Partial differential equations and Hilbert space methods. Wiley, New York, 1980. ISBN 0-471-04089-4. xv + 270 pp. LCCN QA374 .G97.

Guttman:1980:CRF

[98] R. Guttman, S. Lewis, and J. Rinzel. Control of repetitive firing in squid axon membrane as a model for a neurooscillator. Journal of Physiology, 305:377–395, 1980. CODEN JPHYA7. ISSN 0022-3751.

Herbert:1980:CCD

[99] P. D. N. Herbert and T. J. Crease. Clonal coexistence in Daphnia pulex: Another planktonic paradox. Science, 207:1363–1365, 1980. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Irwin:1980:SDS

[100] M. C. Irwin. Smooth Dynamical Systems, volume 94 of Pure and applied mathematics. Academic Press, 1980. ISBN 0-12-374450-4. x + 259 pp. LCCN QA3 .P8 v.94.

Lepeschkin:1980:CPG

[101] E. Lepeschkin, H. C. Herrlich, S. Rush, J. L. Jones, and R. E. Jones. Cardiac potential gradients between defibrillation electrodes. Medical Instrumentation, 14: 57, 1980. CODEN MLISBY. ISSN 0090-6689.

Odell:1980:QTS

[102] G. M. Odell. Qualitative theory of systems of ordinary differential equations, including phase plane analysis and the use of the hopf bifurcation theorem. In Segel [520], pages 649–727. ISBN 0-521-22925-1. LCCN QH506 .M38.

Sanchez-Palencia:1980:LNP

[103] E. Sanchez-Palencia. Non-homogeneous media and vibration theory. In Lecture Notes in Physics, pages ix + 398. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980.

Winfree:1980:GBT

[104] Arthur T. Winfree. The Geometry of Biological Time, volume 8 of Biomathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. ISBN 0-387-09373-7 (New York). 3-540-09373-7 (Berlin). xiii + 530 pp. LCCN QH527 .W55 1980.

Chesson:1981:EVP

[105] P. L. Chesson and R. R. Warner. Environmental variability promotes coexistence in lottery competitive systems. American Naturalist, 117:923–943, 1981. CODEN AMNTA4. ISSN 0003-0147.

Eshel:1981:KSS

[106] I. Eshel and U. Motro. Kin selection and strong evolutionary stability of mutual help. Theoretical Population Biology, 19:420–433, 1981. CODEN TLP-BAQ. ISSN 0040-5809.

Falconer:1981:IQG

[107] D. S. Falconer. Introduction to Quantitative Genetics. Longman Scientific and Technical, Harlow, Essex, UK, second edition, 1981. ISBN 0-470-20474-5, 0-582-44195-1. viii + 340 pp. LCCN QH452.7 .F34 1986.

Guevara:1981:PLP

[108] M. R. Guevara, L. Glass, and A. Shrier.

Phase locking, period doubling bifurcations and irregular dynamics in periodically stimulated cardiac cells. Science, 214:1350–1353, 1981. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Kawato:1981:TSS

[109] M. Kawato. Transient and steady state phase response curves of limit cycle oscillators. Journal of Mathematical Biology, 12:13–30, 1981. CODEN JMBLAJ. ISSN 0303-6812.

Kohn:1981:PIP

[110] A. F. Kohn, A. Freitas da Rocha, and J. P. Segundo. Presynaptic irregularity and pacemaker inhibition. Biological Cybernetics, 41:5–18, 1981. CODEN BICYAF. ISSN 0340-1200.

Nurse:1981:GAC

[111] P. Nurse. Genetics analysis of the cell cycle. In Glover and Hopwood [523], pages 291–315. ISBN 0-521-23748-3. LCCN QR1 .S7 v.31.

Pringle:1981:SCC

[112] J. R. Pringle and L. H. Hartwell. The Saccharomyces cerevisiae cell cycle. In Strathern et al. [522], pages 97–142. ISBN 0-87969-149-2. LCCN QK623.S23 M643 1982.

Reijnders:1981:PDH

[113] P. J. H. Reijnders, H. E. Drescher, J. L. van Haaften, E. Bogebjerg Hansen, and S. Tougaard. Population dynamics of the harbour seal in the Wadden Sea. In Reijnders and Wolff [521], pages 19– 32. ISBN 90-6191-057-9. LCCN QL713.2 .M364 1982.

Goulden:1982:BSE

[114] C. E. Goulden, L. L. Henry, and A. J. Tessier. Body size, energy reserves, and competitive ability in three species of cladocera. Ecology, 63:1780–1789, 1982. CODEN ECOLAR. ISSN 0012-9658.

Guevara:1982:PLP

[115] M. R. Guevara and L. Glass. Phase locking, period doubling bifurcations, and chaos in a mathematical model of a periodically driven oscillator: A theory for the entrainment of biological oscillators and the generation of cardiac dysrhythmias. Journal of Mathematical Biology, 14:1–23, 1982. CODEN JMBLAJ. ISSN 0303-6812.

Hartman:1982:ODE

[116] Philip Hartman. Ordinary Differential Equations. Birkhäuser Boston Inc., Cambridge, MA, USA, second edition, 1982. ISBN 3-7643-3068-6. xiv + 612 pp. LCCN QA372 .H33 1982.

Heiden:1982:DPD

[117] U. an der Heiden and M. C. Mackey. The dynamics of production and destruction: Analytic insight into complex behavior. Journal of Mathematical Biology, 16:75– 101, 1982. CODEN JMBLAJ. ISSN 0303-6812.

Hoppensteadt:1982:PLB

[118] F. C. Hoppensteadt and J. Keener.

Phase locking of biological clocks. Journal of Mathematical Biology, 15:339– 349, 1982. CODEN JMBLAJ. ISSN 0303-6812.

Kingman:1982:EEL

[119] J. F. C. Kingman. Exchangeability and the evolution of large populations. In Koch and Spizzichino [524], pages 97–112. ISBN 0-444-86403-2. LCCN QA273.A1 I55 1981.

Kingman:1982:GLP

[120] J. F. C. Kingman. On the genealogy of large populations. Journal of Applied Probability, 19A:27–43, 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213548.

McQueen:1982:FDM

[121] D. M. McQueen, C. S. Peskin, and E. L. Yellin. Fluid dynamics of the mitral valve: physiological aspects of a mathematical model. American Journal of Physiology, 242:H1095–H1110, 1982. CODEN AJPHAP. ISSN 0002-9513 (print), 2163-5773 (electronic).

Paloheimo:1982:GMD

[122] J. E. Paloheimo, S. J. Crabtree, and W. D. Taylor. Growth model of Daphnia. Canadian Journal of Fisheries and Aquatic Sciences, 39:598–606, 1982. CO-DEN CJFSDX. ISSN 0706-652X.

Watterson:1982:MSA

[123] G. A. Watterson. Mutant substitutions at linked nucleotide sites. Advances in Applied Probability, 14:206–224, 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

Arnold:1983:GMT

[124] V. I. Arnold. Geometrical Methods in the Theory of Ordinary Differential Equations, volume 250 of Grundlehren der mathematischen Wissenschaften. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. ISBN 0-387-90681-9. x + 334 pp. LCCN QA372 .A6913 1983. Translated by J. Szucs; English translation edited by M. Levi.

Berg:1983:RWB

[125] Howard C. Berg. Random Walks in Biology. Princeton University Press, 1983. ISBN 0-691-00064-6. 152 pp. LCCN QH323.5.B45 1993. US$12.95.

Evans:1983:CPS

[126] T. Evans, E. T. Rosenthal, J. Youngbloom, D. Distel, and T. Hunt. Cyclin: a protein specified by maternal mRNA in sea urchin eggs that is destroyed at each cleavage division. Cell, 33:389–396, 1983. CODEN CELLB5. ISSN 0092-8674.

Glass:1983:BCP

[127] L. Glass, M. R. Guevara, A. Shrier, and R. Perez. Bifurcation and chaos in a periodically stimulated cardiac oscillator. Physica, 7D:89–101, 1983. CO-DEN PHYSAG. ISSN 0031-8914 (print), 1873-1767 (electronic).

Guckenheimer:1983:NOD

[128] J. Guckenheimer and P. J. Holmes. Nonlinear Oscillations, Dynamical Systems and Bifurcations of Vector Fields, volume 42 of Applied mathematical sciences. Springer, New York, 1983. ISBN 0-387-90819-6. xii + 459 pp. LCCN QA867.5 .G8 1993.

Hudson:1983:PNA

[129] R. R. Hudson. Properties of a neutral allele model with intragenic recombination. Theoretical Population Biology, 23:183–201, 1983. CODEN TLP-BAQ. ISSN 0040-5809.

Martinez:1983:BSN

[130] O. Diez Martinez and J. P. Segundo. Behavior of a single neuron in a recurrent excitatory loop. Biological Cybernetics, 47:33–41, 1983. CODEN BICYAF. ISSN 0340-1200.

McQueen:1983:CDP

[131] D. M. McQueen and C. S. Peskin.

Computer-assisted design of pivotingdisc prosthetic mitral valves. Journal of thoracic and cardiovascular surgery, 86: 126–135, 1983. CODEN JTCSAQ. ISSN 0022-5223.

Porter:1983:FRF

[132] K. G. Porter, J. J. D. Orcutt, and J. Gerritsen. Functional response and fitness in a generalist filter feeder, Daphnia magna (Cladocera: Crustacea). Ecology, 64:735–742, 1983. CODEN ECOLAR. ISSN 0012-9658.

Tajima:1983:ERD

[133] F. Tajima. Evolutionary relationships of DNA sequences in finite population. Genetics, 105:437–460, 1983. CODEN GENTAE. ISSN 0016-6731.

Gerhart:1984:CCD

[134] J. Gerhart, M. Wu, and M. W. Kirschner. Cell cycle dynamics of an M-phase-specific cytoplasmic factor in Xenopus laevis oocytes and eggs. Journal of Cell Biology, 98:1247–1255, 1984. CODEN JCLBA3. ISSN 0021-9525.

Glass:1984:DPE

[135] L. Glass and A. T. Winfree. Discontinuities in phase-resetting experiments. American Journal of Physiology, 246: R251–R258, 1984. CODEN AJPHAP. ISSN 0002-9513 (print), 2163-5773 (electronic).

Glass:1984:GBP

[136] L. Glass, M. R. Guevara, J. Bélair, and A. Shrier. Global bifurcations of a periodically forced biological oscillator. Physical Review A, 29:1348–1357, 1984. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-

Hopfield:1984:NNG

[137] J. J. Hopfield. Neural networks with graded responses have collective computational properties like those of twostate neurons. Proceedings of the National Academy of Sciences of the United States of America, 81:3088–3092, 1984. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Keener:1984:GBP

[138] J. P. Keener and L. Glass. Global Bifurcations of a periodically forced nonlinear oscillator. Journal of Mathematical Biology, 21:175–190, 1984. CODEN JM-BLAJ. ISSN 0303-6812.

Mackey:1984:DRI

[139] M. C. Mackey and U. an der Heiden. The dynamics of recurrent inhibition. Journal of Mathematical Biology, 19:211– 225, 1984. CODEN JMBLAJ. ISSN 0303-6812.

McMahon:1984:MRL

[140] T. A. McMahon. Muscles, Reflexes, and Locomotion. Princeton University Press, Princeton, NJ, USA, 1984. ISBN 0-691-08322-3, 0-691-02376-X (paperback). xvi + 331 pp. LCCN QP321 .M338 1984.

Saunders:1984:GNS

[141] I. W. Saunders, S. Tavaré, and G. A.

Watterson. On the genealogy of nested subsamples from a haploid population. Advances in Applied Probability, 16: 471–491, 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

Shmida:1984:CPS

[142] A. Shmida and S. Ellner. Coexistence of plant species with similar niches. Vegetatio, 58:29–55, 1984. CODEN VGTOA4. ISSN 0042-3106.

Tyson:1984:CND

[143] J. J. Tyson. The control of nuclear division in Physarum polycephalum. In Nurse and Streiblova [525], pages 175– 190. ISBN 0-442-30778-0, 0-17-771103-5. LCCN QR73.7 .M53 1984.

Drazin:1985:HS

[144] P. G. Drazin and W. H. Reid. Hydrodynamic Stability. Cambridge University Press, New York, NY, 1985.

Ellner:1985:EGS

[145] S. Ellner. ESS germination strategies in randomly varying environments. Theoretical Population Biology, 28:50–116, 1985. CODEN TLPBAQ. ISSN 0040-5809.

Heiden:1985:SPS

[146] U. an der Heiden. Stochastic properties of simple differential-delay equations. In Meinardus and Nurnberger [527], page 351. ISBN 3-7643-1733-7. LCCN QA371 .D37 1985.

Holt:1985:PDT

[147] R. D. Holt. Population dynamics in two-patch environments; some anomalous consequences of an optimal habitat distribution. Theoretical Population Biology, 28:181–208, 1985. CODEN TLP-BAQ. ISSN 0040-5809.

Huxley:1985:CBD

[148] H. E. Huxley and M. Kress. Crossbridge behaviour during muscle contraction. Journal of Muscle Research Cell Motility, 6:153–161, 1985. CODEN JM-RMD3. ISSN 0142-4319.

Kessler:1985:CCP

[149] J. O. Kessler. Co-operative and concentrative phenomena of swimming microorganisms. Contemporary physics, 26: 147–166, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kessler:1985:HFM

[150] J. O. Kessler. Hydrodynamic focusing of motile algal cells. Nature, 313:218–220, 1985. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

McQueen:1985:CDB

[151] D. M. McQueen and C. S. Peskin.

Computer-assisted design of butterfly bileaflet valves for the mitral position. Scandinavian Journal of Thoracic and Cardiovascular Surgery, 19: 139–148, 1985. CODEN SJTCAO. ISSN 0036-5580.

Meisner:1985:ETA

[152] J. S. Meisner, D. M. McQueen, Y. Ishida, H. O. Vetter, U. Bortolotti, J. A. Strom, R. W. M. Frater, C. S. Peskin, and E. L. Yellin. Effects of timing of atrial systole on LV filling and mitral valve closure: computer and dog studies. American Journal of Physiology, 249: H604–H619, 1985. CODEN AJPHAP. ISSN 0002-9513 (print), 2163-5773 (electronic).

Murdoch:1985:TDT

[153] W. W. Murdoch and E. McCauley.

Three distinct types of dynamic behavior shown by a single planktonic system. Nature, 316:628–630, 1985. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Othmer:1985:MAT

[154] H. G. Othmer. The mathematical aspect of temporal oscillations in reacting systems. In Field and Burger [526], pages 7– 54. ISBN 0-471-89384-6. LCCN QD502 .O83 1985. US$85.00.

Othmer:1985:MSR

[155] H. G. Othmer, P. B. Monk, and P. E. Rapp. A model for signal relay and adaptation in Dictyostelium discoideum. Part II. Analytical and numerical results. Mathematical Biosciences, 77:77– 139, 1985. CODEN MABIAR. ISSN 0025-5564.

Petrofsky:1985:CCM

[156] J. S. Petrofsky and C. A. Phillips. Closed-loop control of movement of skeletal muscle. CRC critical reviews in biomedical engineering, 13:35–96, 1985. CODEN CRBEDR. ISSN 0278-940X.

Sommer:1985:GCB

[157] J. R. Sommer and B. Scherer. Geometry of cell and bundle appositions in cardiac muscle: light microscopy. American Journal of Physiology, 248:H792–H803, 1985. CODEN AJPHAP. ISSN 0002-9513 (print), 2163-5773 (electronic).

Woledge:1985:EAM

[158] R. Woledge, N. Curtin, and E. Homsher. Energetic Aspects of Muscle Contraction. Academic Press, New York, NY, 1985. ISBN 0-12-761580-6. xiii + 359 pp. LCCN QP321 .W58 1985.

Yanagida:1985:SDA

[159] T. Yanagida, T. Arata, and F. Oosawa. Sliding distance of actin filament induced by a myosin cross-bridge during one ATP hydrolysis cycle. Nature, 316:366–369, 1985. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Chen:1986:PGF

[160] P-S. Chen, P. D. Wolf, F. J. Claydon, III, E. G. Dixon, H. J. Vidaillet, Jr., N. D. Danieley, T. C. Pilkington, and R. E. Ideker. The potential gradient field created by epicardial defibrillation electrodes in dogs. Circulation, 74:626–636, 1986. CODEN CIRCAZ. ISSN 0009-7322.

Cooke:1986:MMC

[161] R. Cooke. The mechanism of muscle contraction. CRC Critical Reviews in Biochemistry, 21:53–118, 1986. CODEN CRBCAI. ISSN 0045-6411.

Devaney:1986:ICD

[162] Robert L. Devaney. An Introduction to Chaotic Dynamical Systems. W. Benjamin, New York, 1986. ISBN 0-8053-1601-9. xii + 320 pp. LCCN QA614.8 .D48 1986.

Doedel:1986:ASC

[163] E. Doedel. AUTO: Software for continuation and bifurcation problems in ordinary differential equations. Technical report, California Institute of Technology, 1986.

Ellner:1986:GDP

[164] S. Ellner. Germination dimorphisms and parent-offspring conflict in seed germination. Journal of Theoretical Biology, 123:173–185, 1986. CODEN JTBIAP. ISSN 0022-5193.

Glaser:1986:AVR

[165] R. W. Glaser. Appearance of a “critical voltage” in reversible electric breakdown. Studia Biophysica, 16:77–86, 1986. CODEN STBIBN. ISSN 0081-6337.

Guevara:1986:PRS

[166] M. R. Guevara, A. Shrier, and L. Glass. Phase resetting of spontaneously beating embryonic ventricular heart cell aggregates. American Journal of Physiology, 251:H1298–H1305, 1986. CODEN AJPHAP. ISSN 0002-9513 (print), 2163-5773 (electronic).

Hedrick:1986:GPH

[167] P. W. Hedrick. Genetic polymorphism in heterogeneous environments: a decade later. Annual Review of Ecology and Systematics, 17:535–566, 1986. CODEN ARECBC. ISSN 0066-4162.

Kessler:1986:ICF

[168] J. O. Kessler. Individual and collective fluid dynamics of swimming cells. Journal of Fluid Mechanics, 173:191– 205, 1986. CODEN JFLSA7. ISSN 0022-1120.

Metz:1986:DPS

[169] J. A. J. Metz and O. Diekmann. The Dynamics of Physiologically Structured Populations, volume 68 of Lecture Notes in Biomathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. ISBN 0-387-16786-2 (New York), 3-540-16786-2 (Berlin). xii + 511 pp. LCCN QH352 .D961 1986.

Plonsey:1986:EMM

[170] R. Plonsey and R. C. Barr. Effect of microscopic and macroscopic discontinuities on the response of cardiac tissue to defibrillating (stimulating) currents. Medical & Biological Engineering & Computing, 24:130–136, 1986. CO-DEN MBECDY. ISSN 0140-0118.

Powell:1986:QTR

[171] K. T. Powell, E. G. Derrick, and J. C.

Weaver. A quantitative theory of reversible electrical breakdown of bilayer membranes. Bioelectrochemistry and Bioenergetics, 15:243–255, 1986. CO-DEN BEBEBP. ISSN 0302-4598.

Russell:1986:CFI

[172] P. Russell and P. Nurse. cdc25^{+} functions as an inducer in the mitotic control of fission yeast. Cell, 45:145–153, 1986. CODEN CELLB5. ISSN 0092-8674.

Simanis:1986:CCC

[173] V. Simanis and P. Nurse. The cell cycle control gene cdc2 of fission yeast encodes a protein kinase potentially regulated by phosphorylation. Cell, 45:261–268, 1986. CODEN CELLB5. ISSN 0092-8674.

Swann:1986:PPI

[174] K. Swann and M. Whitaker. The part played by inositol trisphosphate and calcium in the propagation of the fertilization wave in sea urchin eggs. Journal of Cell Biology, 103(6 (part 1)):2333–2342, 1986. CODEN JCLBA3. ISSN 0021-9525.

Woods:1986:RTR

[175] N. M. Woods, K. S. R. Cuthberson, and P. H. Cobbold. Repetitive transient rises in cytoplasmic free calcium in hormone-stimulated hepatocytes. Nature, 319:600–602, 1986. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Zachmanoglou:1986:IPD

[176] E. C. Zachmanoglou and Dale W. Thoe. Introduction to Partial Differential Equations with Applications. Dover, 1986. ISBN 0-486-65251-3. x + 405 pp. LCCN QA377 .Z32 1986. US$9.95.

Bull:1987:EPV

[177] J. J. Bull. Evolution of phenotypic variance. Evolution, 41:313–315, 1987. CO-DEN EVOLAO. ISSN 0014-3820.

Cann:1987:MDH

[178] R. Cann, M. Stoneking, and A. C. Wilson. Mitochondrial DNA and human evolution. Nature, 325:31–36, 1987. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Carvalho:1987:CED

[179] G. R. Carvalho and D. J. Crisp. The clonal ecology of Daphnia magna (crustacea: cladocera) I. Temporal changes in the clonal structure of a natural population. Journal of Animal Ecology, 56: 453–468, 1987. CODEN JAECAP. ISSN 0021-8790.

Derrida:1987:LEP

[180] B. Derrida, K. Mecheri, and J. L. Pichard. Lyapounov exponents of products of random matrices: weak disorder expansion. Journal de Physique, 48:733– 740, 1987. CODEN JOPQAG. ISSN 0302-0738.

Glass:1987:UBC

[181] L. Glass, M. R. Guevara, and A. Shrier. Universal bifurcations and the classification of cardiac arrhythmias. Annals of the New York Academy of Sciences, 504:168–178, 1987. CODEN ANYAA9. ISSN 0077-8923 (print), 1749-6632 (electronic).

Heiden:1987:MFA

[182] U. an der Heiden and M. C. Mackey.

Mixed feedback: A paradigm for regular and irregular oscillations. In Rensing et al. [529], pages 30–46. ISBN 0-387-17765-5. LCCN QP84.6 .T44 1987.

Ikeda:1987:HDC

[183] I. Ikeda and K. Matsumoto. High dimensional chaotic behavior in systems with time-delayed feedback. Physica D, 29:223–235, 1987. CODEN PDNPDT. ISSN 0167-2789 (print), 1872-8022 (electronic).

Jordan:1987:NOD

[184] D. W. Jordan and P. Smith. Nonlinear Ordinary Differential Equations. Clarendon Press, Oxford, second edition, 1987. ISBN 0-19-859657-X, 0-19-859656-1 (paperback). ix + 381 pp. LCCN QA372 .J581 1987.

Krassowska:1987:CFS

[185] W. Krassowska, T. C. Pilkington, and R. E. Ideker. The closed form solution to the periodic core-conductor model using asymptotic analysis. IEEE Transactions on Biomedical Engineering, 34 (519–531), 1987. CODEN IEBEAX. ISSN 0018-9294.

Krebs:1987:IBE

[186] J. R. Krebs and N. B. Davies. An Introduction to Behavioral Ecology. Sinauer Associates, Inc., Sunderland, MA, second edition, 1987. ISBN 0-87893-431-6, 0-87893-428-6 (paperback). ix + 389 pp. LCCN QH371 .K731 1987b.

McCauley:1987:CSP

[187] E. McCauley and W. W. Murdoch. Cyclic and stable populations: Plankton as paradigm. American Naturalist, 129: 97–121, 1987. CODEN AMNTA4. ISSN 0003-0147.

Ripley:1987:SS

[188] Brian D. Ripley. Stochastic Simulation. Wiley series in probability and mathematical statistics. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1987. ISBN 0-471-81884-4. xi + 237 pp. LCCN QA76.9.C65 R571 1987. US$29.95.

Rodhouse:1987:CBC

[189] P. G. Rodhouse and C. M. Roden. Carbon budget for a coastal inlet in relation to an intensive cultivation of suspension feeding bivalve molluscs. Marine Ecology Progress Series, 36:225–236, 1987. CODEN MESEDT. ISSN 0171-8630.

Russell:1987:NRM

[190] P. Russell and P. Nurse. Negative regulation of mitosis by wee1^{+}, a gene encoding a protein kinase homolog. Cell, 49: 559–567, 1987. CODEN CELLB5. ISSN 0092-8674.

Seger:1987:WIB

[191] J. Seger and H. J. Brockmann. What is beg-hedging? Oxford Surveys in Evolutionary Biology, 4:182–211, 1987. CO-DEN OSEBE3. ISSN 0265-072X.

Winfree:1987:WTB

[192] A. T. Winfree. When Time Breaks Down: The Three Dimensional Dynamics of Electrochemical Waves and Cardiac Arrhythmia. Princeton University Press, Princeton, NJ, USA, 1987. ISBN 0-691-08443-2, 0-691-02402-2 (paperback). xiv + 339 pp. LCCN QP84.6 .W56 1987.

Andrews:1988:HFO

[193] B. J. Andrews, R. H. Baxendale, R. Barnett, G. F. Phillips, T. Yamazaki, J. P. Paul, and P. Freeman. Hydrid FES orthosis incorporating closed loop control and sensory feedback. Journal of biomedical engineering, 10:189–195, 1988. CODEN JBIEDR. ISSN 0141-5425.

Bohrer:1988:SME

[194] R. N. Bohrer and W. Lampert. Simultaneous measurement of the effect of food concentration on assimilation and respiration in Daphnia magna straus. Functional Ecology, 2:463–471, 1988. CO-DEN FECOE5. ISSN 0269-8463.

Chernysh:1988:MED

[195] A. M. Chernysh, V. Y. Tabak, and M. S.

Bogushevich. Mechanisms of electrical defibrillation of the heart. Resuscitation, 16:169–178, 1988. CODEN RSUSBS. ISSN 0300-9572.

Frazier:1988:EFR

[196] D. W. Frazier, W. Krassowska, P-S. Chen, P. D. Wolf, E. G. Dixon, W. M. Smith, and R. E. Ideker. Extracellular field required for excitation in three-dimensional anisotropic canine myocardium. Circulation Research, 63: 147–164, 1988. CODEN CIRUAL. ISSN 0009-7330.

Glaser:1988:REB

[197] R. W. Glaser, S. L. Leikin, L. V. Chernomordik, V. F. Pastushenko, and A. I.

Sokirko. Reversible electrical breakdown of lipid bilayers: formation and evolution of pores. Biochimica et Biophysica Acta, 940:275–287, 1988. CODEN ABB-PAP. ISSN 0001-5253.

Glass:1988:CCR

[198] L. Glass and M. C. Mackey. From Clocks to Chaos: The Rhythms of Life. Princeton University Press, Princeton, NJ, USA, 1988. ISBN 0-691-08495-5 (hardcover), 0-691-08496-3 (paperback). xvii + 248 pp. LCCN QH527 .G5951 1988.

Guevara:1988:PLR

[199] M. R. Guevara, A. Shrier, and L. Glass. Phase locked rhythms in periodically stimulated heart cell aggregates. American Journal of Physiology, 254:H1–H10, 1988. CODEN AJPHAP. ISSN 0002-9513 (print), 2163-5773 (electronic). Heart Circ. Physiol. 23.

Hairston:1988:IVS

[200] N. G. Hairston, Jr. Interannual variation in seasonal predation: its origin and ecological importance. Limnology and Oceanography, 33:1245–1253, 1988. CO-DEN LIOCAH. ISSN 0024-3590.

Hounsgaard:1988:BDC

[201] J. Hounsgaard, H. Hultborn, B. Jesperson, and O. Kiehn. Bistability of α-motoneurones in the decerebrate cat and in the acute spinal cat after intravenous 5-hydydroxytryptophan. Journal of Physiology, 405:345–367, 1988. CO-DEN JPHYA7. ISSN 0022-3751.

Claydon:1988:VCM

[202] F. J. Claydon, III, T. C. Pilkington, A. S. L. Tang, M. N. Morrow, and R. E. Ideker. A volume conductor model of the thorax for the study of defibrillation fields. IEEE Transactions on Biomedical Engineering, BME-35:331–341, 1988. CODEN IEBEAX. ISSN 0018-9294.

Jacob:1988:RSC

[203] R. Jacob, J. E. Merritt, T. J. Hallam, and T. J. Rink. Repetitive spikes in cytoplasmic calcium evoked by histamine in human endothelial cells. Nature, 335: 40–45, September 1988. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Karlin:1988:NPM

[204] S. Karlin. Non-gaussian phenotypic models of quantitative traits. In Eisen et al. [530], pages 123–144. ISBN 0-87893-900-8, 0-87893-901-6 (paperback).

Koshland:1988:CMS

[205] D. E. Koshland. Chemotaxis as a model second-messenger system. Biochemistry, 27(16):5829–5834, 1988. CO-DEN BICHAW. ISSN 0006-2960.

Kothiyal:1988:TCM

[206] K. P. Kothiyal, B. Shankar, L. J. Fogelson, and N. V. Thakor. Threedimensional computer model of electric fields in internal defibrillation. Proceedings of the IEEE, 76:720–730, 1988. CO-DEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Lin:1988:MAD

[207] C. C. Lin and L. A. Segel. Mathematics Applied to Deterministic Problems in the Natural Sciences. SIAM Press, Philadelphia, PA, 1988. ISBN 0-89871-229-7. xxi + 609 pp. LCCN QA37.2 .L55 1988.

Lohka:1988:PMF

[208] M. J. Lohka, M. K. Hayes, and J. L. Maller. Purification of maturationpromoting factor, an intracellular regulator of early mitotic events. Proceedings of the National Academy of Sciences of the United States of America, 85:3009–3013, 1988. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Longtin:1988:COH

[209] A. Longtin and J. G. Milton. Complex oscillations in the human light reflex with ‘mixed’ and delayed feedback. Mathematical Biosciences, 90:183–199, 1988. CODEN MABIAR. ISSN 0025-5564.

Lupu-Meiri:1988:HAR

[210] M. Lupu-Meiri, H. Shapira, and Y. Oron. Hemispheric asymmetry of rapid chloride responses to inositol trisphosphate and calcium in Xenopus oocytes. FEBS Letters, 240:83–87, 1988. CODEN FEBLAL. ISSN 0014-5793.

McCauley:1988:SMV

[211] E. McCauley, W. W. Murdoch, and S. Watson. Simple models and variation in plankton densities among lakes. American Naturalist, 132:383–403, 1988. CODEN AMNTA4. ISSN 0003-0147.

Milton:1988:IPC

[212] J. G. Milton, A. Longtin, T. H. Kirkham, and G. S. Francis. Irregular pupil cycling as a characteristic abnormality in patients with demyelinative optic neuropathy. American Journal of Ophthalmology, 105:402–407, 1988. CO-DEN AJOPAA. ISSN 0002-9394.

Osterhaus:1988:IVC

[213] A. D. M. E. Osterhaus and E. J. Vedder. Identification of virus causing recent seal deaths. Nature, 338:20, 1988. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Pedley:1988:GBP

[214] T. J. Pedley, N. A. Hill, and J. O.

Kessler. The growth of bioconvection patterns in a uniform suspension of gyrotactic micro-organisms. Journal of Fluid Mechanics, 195:223–237, 1988. CODEN JFLSA7. ISSN 0022-1120.

Reulen:1988:SRD

[215] J. P. H. Reulen, J. T. Marcus, M. J. van Gilst, D. Koops, J. E. Bos, G. Tiesinga, F. R. de Vries, and K. Boshuizen. Stimulation and recording of dynamic pupillary reflex: the IRIS technique. Medical & Biological Engineering & Computing, 26:27–32, 1988. CODEN MBECDY. ISSN 0140-0118.

Schoenberg:1988:CMA

[216] M. Schoenberg. Characterization of the myosin adenosine triphosphate (M•ATP) crossbridge in rabbit and frog skeletal muscle fibers. Biophysical Journal, 54:135–148, 1988. CODEN BIO-JAU. ISSN 0006-3495.

Turelli:1988:PGM

[217] M. Turelli. Population genetic models for polygenic variation and evolution. In Eisen et al. [530], pages 601– 618. ISBN 0-87893-900-8, 0-87893-901-6 (paperback).

Urabe:1988:EFC

[218] J. Urabe. Effect of food conditions on the net production of Daphnia galeata: separate effects on growth and reproduction. Bulletin of Plankton Society of Japan, 35:159–174, 1988. CODEN PSJBAZ. ISSN 0387-8961.

Alberts:1989:MBC

[219] B. Alberts, D. Bray, J. Lewis, M. Raff, K. Roberts, and J. D. Watson. Molecular biology of the cell. Garland Publishing, Inc., New York, NY, USA, second edition, 1989. ISBN 0-8240-3695-6, 0-8240-3696-4 (paperback). xliii + 1294 + 67 pp. LCCN QH581.2 .M64 1994.

Arditi:1989:CPD

[220] R. Arditi and L. R. Ginzburg. Coupling in predator-prey dynamics: Ratiodependence. Journal of Theoretical Biology, 139:311–326, 1989. CODEN JT-BIAP. ISSN 0022-5193.

Barton:1989:EQG

[221] N. H. Barton and M. Turelli. Evolutionary quantitative genetics: How little do we know? Annual Review of Genetics, 23:337–370, 1989. CODEN ARVGB7. ISSN 0066-4197.

Bell:1989:SPM

[222] John B. Bell, Phillip Colella, and Harland M. Glaz. A second-order projection method for the incompressible Navier– Stokes equations. Journal of Computational Physics, 85(2):257–283, December 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

Caswell:1989:MPM

[223] Hal Caswell. Matrix Population Models: Construction, Analysis, and Interpretation. Sinauer Associates, Inc., Sunderland, MA, 1989. ISBN 0-87893-094-9 (hardcover), 0-87893-093-0 (paperback). xiv + 328 pp. LCCN QH352 .C3751 1989. US$28.95 (paperback), US$50.00 (hardcover).

Chernomordik:1989:EBL

[224] L. V. Chernomordik and Y. A. Chizmadzhev. Electrical breakdown of lipid bilayer membranes. phenomenology and mechanism. In Neumann et al. [531], pages 83–95. ISBN 0-306-43043-6. LCCN QH585.5.E48 E44 1989.

Chesson:1989:ISB

[225] P. L. Chesson and S. Ellner. Invasibility and stochastic boundedness in monotonic competition models. Journal of Mathematical Biology, 27:117–138, 1989. CODEN JMBLAJ. ISSN 0303-6812.

Donkin:1989:QSR

[226] P. Donkin, J. Widdows, S. V. Evans, C. M. Worral, and M. Carr. Quantitative structure-activity relationships for the effect of hydrophobic organic chemicals on rate of feeding by mussels (Mytilus edulis). Q. Aquatic Toxicology, 15:277–294, 1989.

Falconer:1989:IQG

[227] D. S. Falconer. Introduction to Quantitative Genetics. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, third edition, 1989. ISBN 0-470-21162-8 (Wiley), 0-582-01642-8 (pbk). xii + 438 pp. LCCN QH452.7 .F341 1989.

Griffiths:1989:GPI

[228] R. C. Griffiths. Genealogical-tree probabilities in the infinitely-many-sites model. Journal of Mathematical Biology, 27:667–68, 1989. CODEN JMBLAJ. ISSN 0303-6812.

Harris:1989:CSC

[229] Elizabeth H. Harris. The Chlamydomonas Sourcebook: a comprehensive guide to biology and laboratory use. Academic Press, San Diego, 1989. ISBN 0-12-326880-X. xiv + 780 pp. LCCN QK569.C486H37 1989.

Hill:1989:GBP

[230] N. A. Hill, T. J. Pedley, and J. O. Kessler. Growth of bioconvection patterns in a suspension of gyrotactic micro-organisms in a layer of finite depth. Journal of Fluid Mechanics, 208: 509–543, 1989. CODEN JFLSA7. ISSN 0022-1120.

Hoyt:1989:DTS

[231] R. H. Hoyt, M. L. Cohen, and J. E. Saffitz. Distribution and three-dimensional structure of intercellular junctions in canine myocardium. Circulation Research, 64:563–574, 1989. CODEN CIRUAL. ISSN 0009-7330.

Jackson:1989:PND

[232] E. A. Jackson. Perspectives of Nonlinear Dynamics. Cambridge University Press, Cambridge, 1989. ISBN 0-521-42632-4 (vol 1, paperback), 0-521-34504-9 (vol. 1), 0-521-42633-2 (vol. 2, paperback), 0-521-35458-7 (vol. 2). LCCN QA845 .J28 1989 v.1-2 (1989-1990). Two volumes.

Lampert:1989:ASV

[233] W. Lampert. The adaptive significance of vertical migration of zooplankton. Functional Ecology, 3:21–27, 1989. CO-DEN FECOE5. ISSN 0269-8463.

Leibold:1989:REE

[234] M. A. Leibold. Resource edibility and the effects of predators and productivity on the outcome of trophic interactions. American Naturalist, 134:922–949, 1989. CODEN AMNTA4. ISSN 0003-0147.

Longtin:1989:ITF

[235] A. Longtin and J. G. Milton. Insight into the transfer function, gain and oscillation onset for the pupil light reflex using delay-differential equations. Biological Cybernetics, 61:51–58, 1989. CO-DEN BICYAF. ISSN 0340-1200.

Longtin:1989:MAO

[236] A. Longtin and J. G. Milton. Modelling autonomous oscillations in the human pupil light reflex using nonlinear delaydifferential equations. Bulletin of Mathematical Biology, 51:605–624, 1989. CO-DEN BMTBAP. ISSN 0092-8240.

Milton:1989:CDB

[237] J. G. Milton, A. Longtin, A. Beuter, M. C. Mackey, and L. Glass. Complex dynamics and bifurcations in neurology. Journal of Theoretical Biology, 138:129– 147, 1989. CODEN JTBIAP. ISSN 0022-5193.

Monk:1989:CAO

[238] P. B. Monk and H. G. Othmer. Cyclic AMP oscillations in suspensions of Dictyostelium discoideum. Philosophical transactions of the Royal Society of London, 323(1215):185–224, 1989. CODEN PTRSAV. ISSN 0370-2316.

Murray:1989:MB

[239] J. D. Murray. Mathematical Biology, volume 19 of Biomathematics. Springer, New York, 1989. ISBN 0-387-19460-6 (New York), 3-540-19460-6 (Berlin). xiv + 767 pp. LCCN QH323.5 .M881 1989.

Nisbet:1989:SPM

[240] R. M. Nisbet, W. S. C. Gurney, W. W. Murdoch, and E. McCauley. Structured population models: A tool for linking effects at individual and population level. Biological Journal of the Linnean Society, 37:79–99, 1989. CODEN BJLSBG. ISSN 0024-4066.

Pate:1989:MCA

[241] E. Pate and R. Cooke. A model of crossbridge action: the effects of ATP, ADP, and pi. Journal of Muscle Research Cell Motility, 10:181–196, 1989. CODEN JMRMD3. ISSN 0142-4319.

Pease:1989:MPG

[242] C. M. Pease, R. Lande, and J. J. Bull. A model of population growth, dispersal, and evolution in a changing environment. Ecology, 70:1657–1664, 1989. CODEN ECOLAR. ISSN 0012-9658.

Peskin:1989:FLV

[243] C. S. Peskin. Fiber-architecture of the left ventricular wall: an asymptotic analysis. Communications on Pure and Applied Mathematics (New York), 42(1):79–113, January 1989. CODEN CPAMAT, CPMAMV. ISSN 0010-3640 (print), 1097-0312 (electronic).

Phillipi:1989:HOE

[244] T. Phillipi and J. Seger. Hedging one’s evolutionary bets, revisited. Trends in Ecology and Evolution, 4:41–44, 1989. CODEN TREEEQ. ISSN 0169-5347.

Rooney:1989:CCC

[245] Thomas A. Rooney, Ellen J. Sass, and Andrew P. Thomas. Characterization of cytosolic calcium oscillations induced by phenylephrine and vasopressin in single fura-2-loaded hepatocytes. Journal of biological chemistry, 264(29): 17131–17141, October 1989. CODEN JBCHA3. ISSN 0021-9258.

Stasio:1989:SBF

[246] B. T. De Stasio. The seed bank of a freshwater crustacean: copepodology for the plant ecologist. Ecology, 70:1377– 1389, 1989. CODEN ECOLAR. ISSN 0012-9658.

Taylor:1989:ESO

[247] P. B. Taylor. Evolutionary stability in one-parameter models under weak selection. Theoretical Population Biology, 36:125–143, 1989. CODEN TLPBAQ. ISSN 0040-5809.

Zauderer:1989:PDE

[248] Erich Zauderer. Partial differential equations of applied mathematics. John Wiley and Sons, New York, second edition, 1989. ISBN 0-471-61298-7. xiii + 891 pp. LCCN QA377 .Z38 1988.

Acheson:1990:EFD

[249] D. J. Acheson. Elementary Fluid Dynamics. Oxford applied mathematics and computing science series. Oxford University Press, 1990. ISBN 0-19-859660-X (case), 0-19-859679-0 (paperback). ix + 397 pp. LCCN TA357.A276 1990. US$67.50 (case), US$22.50 (paperback).

Alexander:1990:RSF

[250] J. C. Alexander, E. J. Doedel, and H. G. Othmer. On the resonance structure in a forced excitable system. SIAM Journal on Applied Mathematics, 50(5):1373–1418, October 1990. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Amann:1990:ODE

[251] Herbert Amann. Ordinary Differential Equations: An introduction to nonlinear analysis. de Gruyter, Berlin, 1990. ISBN 0-89925-552-3, 3-11-011515-8. xii + 458 pp. LCCN QA372 .A4313 1990. US$96.00. Translated from the German by Gerhard Metzen.

Arrowsmith:1990:IDS

[252] D. K. Arrowsmith and C. M. Place. An Introduction to Dynamical Systems. Cambridge University Press, 1990. ISBN 0-521-31650-2 (paperback), 0-521-30362-1 (hardcover). 423 pp. LCCN QA614.8 .A77 1990.

Berridge:1990:ICM

[253] M. J. Berridge. Inositol 1,4,5trisphosphate-induced calcium mobilization is localized in Xenopus oocytes. Proceedings of the Royal Society of London. Series B. Biological sciences, 238: 235–343, 1990. CODEN PRLBA4. ISSN 0962-8452 (print), 1471-2954 (electronic).

Bressloff:1990:NDB

[254] P. C. Bressloff and J. Stark. Neuronal dynamics based on discontinuous circle maps. Physics Letters A, 150:187–195, 1990. CODEN PYLAAG. ISSN 0375-9601 (print), 1873-2429 (electronic).

DeLisle:1990:EIT

[255] S. DeLisle, K. H. Krause, G. Denning, B. V. L. Potter, and M. J. Welsh. Effect of inositol trisphosphate and calcium on oscillating elevations of intracellular calcium in Xenopus oocytes. Journal of biological chemistry, 265:11726– 11730, 1990. CODEN JBCHA3. ISSN 0021-9258.

Frank:1990:EVE

[256] S. A. Frank and M. Slatkin. Evolution in a variable environment. American Naturalist, 136:244–260, 1990. CODEN AM-NTA4. ISSN 0003-0147.

Gilmour:1990:PRC

[257] R. F. Gilmour. Phase resetting of circus movement reentry in cardiac tissue. In Zipes and Jalife [532], pages 396–402. ISBN 0-7216-1408-6. LCCN WG330 C38 1990.

Goldbeter:1990:MMS

[258] A. Goldbeter, G. Dupont, and M. J.

Berridge. Minimal model for signalinduced Ca oscillations and for their frequency encoding through protein phosphorylation. Proceedings of the National Academy of Sciences of the United States of America, 87:1461–1465, 1990. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Gurney:1990:PED

[259] W. S. C. Gurney, E. McCauley, R. M.

Nisbet, and W. W. Murdoch. The physiological ecology of Daphnia: A dynamic model of growth and reproduction. Ecology, 71(2):716–732, April 1990. CODEN ECOLAR. ISSN 0012-9658.

Hairston:1990:FSR

[260] N. G. Hairston, Jr. and T. A. Dillon.

Fluctuating selection and response in a population of freshwater copepods. Evolution, 44:1796–1805, 1990. CODEN EVOLAO. ISSN 0014-3820.

Harada:1990:MCA

[261] Y. Harada, T. Sakurada, T. Aoki, D. D. Thomas, and T. Yanagida. Mechanochemical coupling in actomyosin energy transduction studies in an in vitro movement assay. Journal of Molecular Biology, 216:49–68, 1990. CO-DEN JMOBAK. ISSN 0022-2836.

Heiden:1990:MOB

[262] U. an der Heiden and K. Reichard. Multitude of oscillatory behavior in a nonlinear second order differential-difference equation. Zeitschrift für Angewandte Mathematik und Mechanik, 70:621–624, 1990. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Heiden:1990:OMN

[263] U. an der Heiden, A. Longtin, M. C. Mackey, J. G. Milton, and R. Scholl. Oscillatory modes in a nonlinear secondorder differential equation with delay. Journal of Dynamics and Differential Equations, 2(4):423–449, October 1990. CODEN JDDEEH. ISSN 1040-7294.

Kleinfeld:1990:CCI

[264] D. Kleinfeld, F. Raccuia-Behling, and H. J. Chiel. Circuits constructed from identified Aplysia neurons exhibit multiple patterns of persistent activity. Biophysical Journal, 57(4):697–715, 1990. CODEN BIOJAU. ISSN 0006-3495.

Krassowska:1990:PDT

[265] W. Krassowska, D. W. Frazier, T. C. Pilkington, and R. E. Ideker. Potential distribution in three-dimensional periodic myocardium: Part II. application to extracellular stimulation. IEEE Transactions on Biomedical Engineering, BME-37(3):267–284, March 1990. CODEN IEBEAX. ISSN 0018-9294.

Longtin:1990:NCB

[266] A. Longtin, J. G. Milton, J. E. Bos, and M. C. Mackey. Noise and critical behavior of the pupil light reflex at oscillation onset. Physical Review A, 41:6992–7005, 1990. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

McCauley:1990:GRM

[267] E. McCauley, W. W. Murdoch, and R. M. Nisbet. Growth, reproduction, and mortality of Daphnia pulex: life at low food. Functional Ecology, 4:505–514, 1990. CODEN FECOE5. ISSN 0269-8463.

McCauley:1990:PDE

[268] E. McCauley and W. W. Murdoch.

Predator-prey dynamics in environments rich and poor in nutrients. Nature, 343(6257):455–457, February 1, 1990. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Milton:1990:CDN

[269] J. G. Milton, U. an der Heiden, A. Longtin, and M. C. Mackey. Complex dynamics and noise in simple neural networks with delayed mixed feedback. Biomedica Biochimica Acta, 49: 697–707, 1990. CODEN BBIADT. ISSN 0232-766X.

Milton:1990:EPC

[270] John G. Milton and Andre Longtin.

Evaluation of pupil constriction and dilation from cycling measurements. Vision Research (Oxford), 30(4):515–526, 1990. CODEN VISRAM. ISSN 0042-

Nurse:1990:UCM

[271] P. Nurse. Universal control mechanism regulating onset of M-phase. Nature, 344:503–508, 1990. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Ott:1990:CC

[272] E. Ott, C. Grebogi, and J. A. Yorke. Controlling chaos. Physical Review Letters, 64:1196–1199, 1990. CODEN PRL-TAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Pedley:1990:NCM

[273] T. J. Pedley and J. O. Kessler. A new continuum model for suspensions of gyrotactic micro-organisms. Journal of Fluid Mechanics, 212:155–182, March 1990. CODEN JFLSA7. ISSN 0022-1120.

Quan:1990:UBR

[274] W. Quan and Y. Rudy. Unidirectional block and reentry of cardiac excitation: A model study. Circulation Research, 66(2):367–382, February 1990. CODEN CIRUAL. ISSN 0009-7330.

Sepulveda:1990:FEA

[275] N. G. Sepulveda, J. P. Wikswo, Jr., and D. S. Echt. Finite element analysis of cardiac defibrillation current distributions. IEEE Transactions on Biomedical Engineering, BME-37(4):354–365, April 1990. CODEN IEBEAX. ISSN 0018-9294.

Solomon:1990:CAP

[276] M. J. Solomon, M. Glotzer, T. H. Lee, M. Philippe, and M. W. Kirschner. Cyclin activation of p34cdc2. Cell, 63:

1013–1024, 1990. CODEN CELLB5. ISSN 0092-8674.

Tett:1990:LLS

[277] P. Tett. Chapter 4: The photic zone.

In P. J. Henning, A. K. Campbell, M. Whitfield, and L. Maddock, editors, Light and Life in the Sea. Cambridge University Press, New York, NY, 1990.

Tuljapurkar:1990:PDV

[278] Shripad Tuljapurkar. Population Dynamics in Variable Environments, volume 85 of Lecture Notes in Biomathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 0-387-52482-7. 154 pp. LCCN HB849.51 .T85 1990.

Uyeda:1990:MSS

[279] T. Q. P. Uyeda, S. J. Kron, E. M. Mc-Nalley, K. R. Niebling, C. Toyoshima, and J. A. Spudich. The myosin step size: estimation from slow sliding movement of actin over low densities of heavy meromyosin. Journal of Molecular Biology, 214:699–710, 1990. CODEN JMOBAK. ISSN 0022-2836.

Volpe:1990:RIC

[280] P. Volpe, B. H. Alderson-Lang, and G. A. Nickols. Regulation of inositol 1,4,5-trisphosphate-induced Ca^{2+} release. I. effect of Mg^{2+}. American Journal of Physiology, 258(6):C1077– C1085, June 1990. CODEN AJPHAP. ISSN 0002-9513 (print), 2163-5773 (electronic).

Wilder:1990:VNS

[281] B. J. Wilder. Vagus nerve stimulation for the control of epilepsy. Epilepsia, 1990. CODEN EPILAK. ISSN 0013-9580. Supplement 2.

Zeng:1990:CRR

[282] W. Zeng, J. Morissette, R. Brochu, L. Glass, and A. Shrier. Complex rhythms resulting from overdrive suppression in electrically stimulated heart cell aggregates. Pacing and Clinical Electrophysiology: PACE, 13(12):1678– 1685, December 1990. ISSN 0147-8389.

Zeng:1990:TCP

[283] W. Zeng, M. Courtemanche, L. Sehn, A. Shrier, and L. Glass. Theoretical computation of phase locking in embryonic atrial heart cell aggregates. Journal of Theoretical Biology, 145:225–244, 1990. CODEN JTBIAP. ISSN 0022-5193.

Bezprozvanny:1991:BCC

[284] I. Bezprozvanny, J. Watras, and B. E. Ehrlich. Bell-shaped calcium-response curves of Ins(1,4,5)P_{3}- and calciumgated channels from endoplasmic recticulum of cerebellum. Nature, 351: 751–754, 1991. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Bradshaw:1991:CLG

[285] A. D. Bradshaw. The Croonian Lecture 1991: Genostasis and the limits to evolution. Philosophical Transactions of the Royal Society of London Series B, Biological sciences, 333:289–305, 1991.

Cobbold:1991:HCO

[286] P. H. Cobbold, A. Sanchez-Bueno, and C. J. Dixon. The hepatocyte calcium oscillator. Cell calcium, 12:87–95, 1991. CODEN CECADV. ISSN 0143-4160.

Delisle:1991:FDC

[287] S. Delisle. The four dimensions of calcium signalling in Xenopus oocytes. Cell calcium, 12:217–227, 1991. CODEN CECADV. ISSN 0143-4160.

Dillon:1991:ORR

[288] S. M. Dillon. Optical recordings in the rabbit heart show that defibrillation strength shocks prolong the duration of depolarization and the refractory period. Circulation Research, 69:842–856, 1991. CODEN CIRUAL. ISSN 0009-7330.

Dupont:1991:SOP

[289] G. Dupont, M. J. Berridge, and A. Goldbeter. Signal-induced Ca oscillations: Properties of a model based on Cainduced Ca release. Cell calcium, 12: 73–85, 1991. CODEN CECADV. ISSN 0143-4160.

Finch:1991:CCI

[290] E. A. Finch, T. J. Turner, and S. M.

Goldin. Calcium as a coagonist of inositol 1,4,5-trisphosphate-induced calcium release. Science, 252:443–446, 1991. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Gatto:1991:SRM

[291] M. Gatto. Some remarks on models of plankton densities in lakes. American Naturalist, 137:264–267, 1991. CODEN AMNTA4. ISSN 0003-0147.

Gillespie:1991:CME

[292] J. H. Gillespie. The Causes of Molecular Evolution. Oxford University Press, Oxford, UK, 1991. ISBN 0-19-506883-1, 0-19-509271-6 (paperback). xiv + 336 pp. LCCN QH455 .G56 1991.

Gilpin:1991:MDE

[293] M. Gilpin and I. Hanski, editors. Metapopulation Dynamics: Empirical and Theoretical Investigations, volume 42 of Biological Journal of the Linnean Society. Academic Press, New York, NY, 1991. ISBN 0-12-284120-4. 336 pp. LCCN ????

Hale:1991:DB

[294] Jack K. Hale and Huseyin Kocak. Dynamics and Bifurcations. Springer-Verlag, New York, 1991. ISBN 0-387-97141-6, 3-540-97141-6. xiv + 568 pp. LCCN QA372 .H183 1991.

Hanski:1991:SMD

[295] I. Hanski. Single-species metapopulation dynamics: Concepts, models, and observations. In Gilpin and Hanski [293], pages 89–103. ISBN 0-12-284120-4. LCCN ????

Harootunian:1991:GCO

[296] A. T. Harootunian, J. P. Y. Kao, S. Paranjape, and R. Y. Tsien. Generation of calcium oscillations in fibroblasts by positive feedback between calcium and IP3. Science, 251:75–78, 1991. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Higuchi:1991:SDB

[297] H. Higuchi and Y. E. Goldman. Sliding distance between actin and myosin filaments per ATP molecule hydrolysed in skinned muscle fibres. Nature, 352 (6333):352–354, July 25, 1991. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Hudson:1991:GGC

[298] R. R. Hudson. Gene genealogies and the coalescent process. Oxford Surveys in Evolutionary Biology, 7:1–44, 1991. CO-DEN OSEBE3. ISSN 0265-072X.

Lechleiter:1991:SCW

[299] J. Lechleiter, S. Girard, E. Peralta, and D. Clapham. Spiral calcium wave propagation and annihilation in xenopus laevis oocytes. Science, 252:123– 126, 1991. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Ludwig:1991:ESP

[300] D. Ludwig and S. A. Levin. Evolutionary stability of plant communities and the maintenance of multiple dispersal types. Theoretical Population Biology, 40:285–307, 1991. CODEN TLPBAQ. ISSN 0040-5809.

Luo:1991:MVC

[301] C-H. Luo and Y. Rudy. A model of the ventricular cardiac action potential: Depolarization, repolarization, and their interaction. Circulation Research, 68: 1501–1526, 1991. CODEN CIRUAL. ISSN 0009-7330.

Malcolm:1991:MPB

[302] I. Malcolm. Motor proteins: Biomechanics goes quantum. Nature, 352(6333): 284–286, July 25, 1991. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Metz:1991:EFD

[303] J. A. J. Metz and O. Diekmann. Exact finite dimensional representations of models for physiologically structured populations. I: The abstract foundations of linear chain trickery. In Goldstein et al. [534], pages 269–289. ISBN 0-8247-8571-1. LCCN QA370 .D56 1991.

Meyer:1991:CS

[304] Tobias Meyer and Lubert Stryer. Calcium spiking. Annual review of biophysics and biophysical chemistry, 20: 153–174, 1991. CODEN ARBCEY.

Nisbet:1991:PDE

[305] R. M. Nisbet, E. McCauley, A. M. de Roos, W. W. Murdoch, and W. S. C. Gurney. Population dynamics and element recycling in an aquatic plantherbivore system. Theoretical Population Biology, 40(2):125–147, October 1991. CODEN TLPBAQ. ISSN 0040-5809.

Nuccitelli:1991:HDS

[306] R. Nuccitelli. How do sperm activate eggs? Current topics in developmental biology, 25:1–16, 1991. CODEN CT-DBA5. ISSN 0070-2153.

Othmer:1991:DFE

[307] H. G. Othmer. The dynamics of forced excitable systems. In Holden et al. [536], pages 213–232. ISBN 0-306-43800-3. LCCN QA927 .N27 1989.

Perko:1991:DED

[308] L. Perko. Differential Equations and Dynamical Systems, volume 7 of Texts in applied mathematics. Springer, New York, 1991. ISBN 0-387-97443-1 (New York), 3-540-97443-1 (Berlin). xii + 403 pp. LCCN QA372 .P47 1991.

Renshaw:1991:MBP

[309] E. Renshaw. Modelling Biological Populations in Space and Time. Cambridge University Press, Cambridge, UK, 1991.

ISBN 0-521-30388-5 (hardback), 0-521-44855-7 (paperback). xvii + 403 pp. LCCN QH352 .R46 1991.

Rooney:1991:OCC

[310] T. A. Rooney, D. C. Renard, E. J. Sass, and A. P. Thomas. Oscillatory cytosolic calcium waves independent of stimulated inositol 1,4,5-trisphosphate formation in hepatocytes. Journal of biological chemistry, 266(19):12272–12282, 1991. CODEN JBCHA3. ISSN 0021-9258.

Somogyi:1991:HCO

[311] R. Somogyi and J. W. Stucki. Hormoneinduced calcium oscillations in liver cells can be explained by a simple one pool model. Journal of biological chemistry, 266(17):11068–11077, 1991. CODEN JBCHA3. ISSN 0021-9258.

Takada:1991:DAE

[312] T. Takada and J. Kigami. The dynamical attainability of ESS in evolutionary games. Journal of Mathematical Biology, 29:513–529, 1991. CODEN JMBLAJ. ISSN 0303-6812.

Taylor:1991:KSA

[313] Edwin W. Taylor. Kinetic studies on the association and dissociation of myosin subfragment 1 and actin. Journal of biological chemistry, 266(1):294–302, January 5, 1991. CODEN JBCHA3. ISSN 0021-9258.

Thomas:1991:STO

[314] A. P. Thomas, D. C. Renard, and T. A.

Rooney. Spatial and temporal organization of calcium signalling in hepatocytes. Cell calcium, 12:111–126, 1991. CODEN CECADV. ISSN 0143-4160.

Tuljapurkar:1991:PDV

[315] S. Tuljapurkar. Population Dynamics in Variable Environments, volume 85 of Lecture Notes in Biomathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-52482-7. 154 pp. LCCN HB849.51 .T85 1990.

Urabe:1991:EFC

[316] J. Urabe and Y. Watanabe. Effect of food concentration on the assimilation and production efficiencies of Daphnia galeata. Functional Ecology, 5:635–641, 1991. CODEN FECOE5. ISSN 0269-8463.

Watras:1991:ICC

[317] J. Watras, I. Bezprozvanny, and B. E. Ehrlich. Inositol 1,4,5-trisphosphate-gated channels in cerebellum: Presence of multiple conductance states. Journal of Neuroscience, 11(10):3239–3245, October 1991. CODEN JNRSDS. ISSN 0270-6474.

Yarom:1991:RHS

[318] Y. Yarom. Rhythmogenesis in a hybrid system interconnecting an olivary neuron to an analog network of coupled oscillators. Neuroscience, 44(2):263–275, 1991. CODEN NRSCDN. ISSN 0306-4522.

Zeng:1991:ERD

[319] W. Zeng, L. Glass, and A. Shrier. Evolution of rhythms during periodic stimulation of embryonic chick heart cell aggregates. Circulation Research, 69(4):1022– 1033, October 1991. CODEN CIRUAL. ISSN 0009-7330.

Aida:1992:OML

[320] T. Aida and P. Davis. Oscillation modes of laser diode pumped hybrid bistable system with large delay and application to dynamic memory. IEEE Journal of Quantum Electronics, 28(3):686–699, March 1992. CODEN IEJQA7. ISSN 0018-9197.

Brown:1992:EHE

[321] J. S. Brown and N. B. Pavlovic. Evolution in heterogeneous environments: Effect of migration on habitat specialization. Evolutionary Ecology, 6:360–382, 1992. CODEN EVECEJ. ISSN 0269-7653.

Garfinkel:1992:CCC

[322] A. Garfinkel, M. Spano, W. L. Ditto, and J. Weiss. Controlling cardiac chaos. Science, 257(5074):1230–1235, August 28, 1992. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Girard:1992:TMC

[323] S. Girard, A. Luckhoff, J. Lechleiter, J. Sneyd, and D. Clapham. Twodimensional model of calcium waves reproduces the patterns observed Xenopus oocytes. Biophysical Journal, 61: 509–517, 1992. CODEN BIOJAU. ISSN 0006-3495.

Glazier:1992:EAR

[324] D. S. Glazier and P. Calow. Energy allocation rules in Daphnia magna — clonal and age differences in the effects of food limitation. Oecologia, 90(4):540– 549, July 1992. CODEN OECOBX. ISSN 0029-8549.

Heide-Jorgensen:1992:ESD

[325] M. P. Heide-Jorgensen and T. Harkonen. Epizootiology of the seal disease in the eastern North Sea. Journal of Applied Ecology, 29:99–107, 1992. CODEN JAPEAI. ISSN 0021-8901.

Holt:1992:AAH

[326] R. D. Holt and M. S. Gaines. Analysis of adaptation in heterogeneous landscapes: Implications for the evolution of fundamental niches. Evolutionary Ecology, 7: 433–447, 1992. CODEN EVECEJ. ISSN 0269-7653.

Hudson:1992:HWG

[327] R. R. Hudson. The how and why of generating gene genealogies. In Takahata and Clark [538], pages 23–36. ISBN 4-7622-6718-X, 0-87893-825-7. LCCN QH352 .M43 1993.

Keizer:1992:TRA

[328] J. Keizer and G. W. De Young. Two roles for Ca in agonist stimulated Ca oscillations. Biophysical Journal, 61:649– 660, 1992. CODEN BIOJAU. ISSN 0006-3495.

Lechleiter:1992:MMI

[329] J. Lechleiter and D. Clapham. Molecular mechanisms of intracellular calcium excitability in X. laevis oocytes. Cell, 69: 283–294, 1992. CODEN CELLB5. ISSN 0092-8674.

McPeek:1992:EDS

[330] M. A. McPeek and R. D. Holt. The evolution of dispersal in spatially and temporally varying environments. American Naturalist, 140:1010–1027, 1992. CO-DEN AMNTA4. ISSN 0003-0147.

Miyazaki:1992:BCW

[331] S. Miyazaki, M. Yuzaki, K. Nakada, H. Shirakawa, S. Nakanishi, S. Nakade, and K. Mikoshiba. Block of Ca^{2+} wave and Ca^{2+} oscillation by antibody to the inositol 1,4,5-trisphosphate receptor in fertilized hamster eggs. Science, 257:251–255, 1992. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Murdoch:1992:IMC

[332] W. W. Murdoch, E. McCauley, R. M.

Nisbet, W. S. C. Gurney, and A. M. de Roos. Individual-based models: Combining testability and generality. In DeAngelis and Gross [537], pages 18– 35. ISBN 0-412-03161-2 (hardcover), 0-412-03171-X (paperback). LCCN QH352 .I53 1992. US$65.00 (hardcover), US$29.00 (paperback).

Nagylaki:1992:ITP

[333] T. Nagylaki. Introduction to Theoretical Population Genetics, volume 21 of Biomathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 3-540-53344-3 (Berlin), 0-387-53344-3 (New York). xi + 369 pp. LCCN QH455 .N32 1992.

Pedley:1992:HPS

[334] T. J. Pedley and J. O. Kessler. Hydrodynamic phenomena in suspensions of swimming micro-organisms. Annual Review of Fluid Mechanics, 24:313–358, 1992. CODEN ARVFA3. ISSN 0066-4189.

Peskin:1992:CFD

[335] C. S. Peskin and D. M. McQueen.

Cardiac fluid dynamics. CRC critical reviews in biomedical engineering, 20(5/6):451–459, 1992. CODEN CRBEDR. ISSN 0278-940X.

Roff:1992:ELH

[336] Derek A. Roff. The Evolution of Life Histories: Theory and Analysis. Chapman and Hall, Ltd., London, UK, 1992. ISBN 0-412-02391-1 (paper), 0-412-02381-4 (cloth). xii + 535 pp. LCCN QH323.5 .R62 1992.

Sharp:1992:AES

[337] A. A. Sharp, L. F. Abbott, and E. Marder. Artificial electrical synapses in oscillatory networks. Journal of Neurophysiology, 67(6):1691–1694, June 1992. CODEN JONEA4. ISSN 0022-3077.

Stearns:1992:ELH

[338] Stephen C. Stearns. The Evolution of Life Histories. Oxford University Press, Oxford, UK, 1992. ISBN 0-19-857741-9. xii + 249 pp. LCCN QH371 .S72 1992.

Tang:1992:TPG

[339] A. S. L. Tang, P. D. Wolf, Y. Afework, W. M. Smith, and R. E. Ideker. Threedimensional potential gradient fields generated by intracardiac catheter and cutaneous patch electrodes. Circulation, 85(5):1857–1864, May 1992. CODEN CIRCAZ. ISSN 0009-7322.

Wharton:1992:CPP

[340] J. M. Wharton, P. D. Wolf, W. M. Smith, P-S. Chen, D. W. Frazier, S. Yabe, N. Danieley, and R. E. Ideker. Cardiac potential and potential gradient fields generated by single, combined, and sequential shocks during ventricular defibrillation. Circulation, 85:1510–1523, 1992. CODEN CIRCAZ. ISSN 0009-7322.

Young:1992:SIM

[341] G. De Young and J. Keizer. A single-pool inositol 1,4,5-trisphophate-receptor-based model for agoniststimulated oscillations in Ca^{+2} concentration. Proceedings of the National Academy of Sciences of the United States of America, 89:9895–9899, 1992. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Zeng:1992:TPR

[342] W. Zeng, L. Glass, and A. Shrier. The topology of phase response curves induced by single and paired stimuli. Journal of Biological Rhythms, 7(2):89–104, Summer 1992. CODEN JBRHEE. ISSN 0748-7304.

Bagshaw:1993:MC

[343] C. R. Bagshaw. Muscle Contraction.

Chapman and Hall, Ltd., London, UK, second edition, 1993. ISBN 0-412-40370-6. x + 155 pp. LCCN QP321 .B26 1993.

Carpenter:1993:TCL

[344] S. R. Carpenter and J. F. Kitchell, editors. The Trophic Cascade in Lakes. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-43145-X. xiv + 385 pp. LCCN QH541.5.L3 T76 1993.

Charnov:1993:LHI

[345] Eric L. Charnov. Life History Invariants: Some explorations of symmetry in evolutionary ecology. Oxford University Press, Oxford, UK, 1993. ISBN 0-19-854072-8 (hardback), 0-19-854071-X (paperback). xv + 167 pp. LCCN QH371 .C485 1993.

Courtemanche:1993:IPP

[346] M. Courtemanche, L. Glass, and J. Keener. Instabilities of a propagating pulse in a ring of excitable media. Physical Review Letters, 70(14):2182– 2185, April 5, 1993. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Hale:1993:IFD

[347] J. K. Hale and S. M. V. Lunel. Introduction to Functional Differential Equations, volume 99 of Applied mathematical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-94076-6 (New York), 3-540-94076-6 (Berlin). x + 447 pp. LCCN QA372 .H35 1993.

Henriquez:1993:SEB

[348] C. S. Henriquez. Simulating the electrical behavior of cardiac tissue using the bidomain model. CRC critical reviews in biomedical engineering, 21:1–77, 1993. CODEN CRBEDR. ISSN 0278-940X.

Holloway:1993:AII

[349] S. L. Holloway, M. Glotzer, R. W. King, and A. W. Murray. Anaphase is initiated by proteolysis rather than by the inactivation of maturation-promoting factor. Cell, 73:1393–1402, 1993. CODEN CELLB5. ISSN 0092-8674.

Josephson:1993:REV

[350] M. E. Josephson, D. Callans, J. M. Almendral, B. G. Hook, and R. B. Kleiman. Resetting and entrainment of ventricular tachycardia associated with infarction: Clinical and experimental studies. In Josephson and Wellens [540], pages 505–536. ISBN 0-87993-549-9. LCCN WG 330 T117-3 1993.

Knisley:1993:OMT

[351] S. B. Knisley, T. F. Blitchington, B. C.

Hill, A. O. Grant, W. M. Smith, T. C. Pilkington, and R. E. Ideker. Optical measurements of transmembrane potential changes during electric field stimulation of ventricular cells. Circulation Research, 72(2):255–270, February 1993. CODEN CIRUAL. ISSN 0009-7330.

Krebs:1993:BEE

[352] J. R. Krebs and N. B. Davies. Behavioral Ecology: An Evolutionary Approach. Blackwell Scientific Publications, Oxford, UK, third edition, 1993. ISBN 0-632-02701-0 (hardback), 0-632-02702-9 (paperback). xi + 482 pp. LCCN QL751 .B345 1991.

Kretzschmar:1993:PMZ

[353] M. Kretzschmar, R. M. Nisbet, and E. McCauley. A predator-prey model for zooplankton grazing on competing algal populations. Theoretical Population Biology, 44(1):32–66, August 1993. CO-DEN TLPBAQ. ISSN 0040-5809.

Lenski:1993:ERE

[354] R. E. Lenski and A. F. Bennett. Evolutionary response of Escherichia coli to thermal stress. American Naturalist, 142:S47–S64, 1993. CODEN AMNTA4. ISSN 0003-0147.

Losson:1993:SMF

[355] J. Losson, M. C. Mackey, and A. Longtin. Solution multistability in first-order nonlinear differential delay equations. Chaos (Woodbury, NY), 3: 167–176, 1993. CODEN CHAOEH. ISSN 1054-1500.

Lynch:1993:EER

[356] M. Lynch and R. Lande. Evolution and extinction in response to environmental change. In Kareiva et al. [541], pages 234–250. ISBN 0-87893-429-4, 0-87893-430-8 (paperback). LCCN QH546 .B55 1993. US$55.00, US$35.00 (paperback).

Murray:1993:CCI

[357] Andrew Murray and Tim Hunt. The cell cycle: an introduction. W. H. Freeman, New York, NY, USA, 1993. ISBN 0-7167-7044-X (hard), 0-7167-7046-6 (soft), 0-19-509529-4. xii + 251 pp. LCCN H605 .M95 1993.

Murray:1993:TM

[358] A. W. Murray. Turning on mitosis. Current biology, 3:291–293, 1993. CODEN CUBLE2. ISSN 0960-9822.

Neu:1993:HST

[359] J. C. Neu and W. Krassowska. Homogenization of syncytial tissues. CRC critical reviews in biomedical engineering, 21:137–199, 1993. CODEN CRBEDR. ISSN 0278-940X.

Novak:1993:NAC

[360] B. Novak and J. J. Tyson. Numerical analysis of a comprehensive model of M-phase control in Xenopus oocyte extracts and intact embryos. Journal of Cell Science, 106:1153–1168, 1993. CO-DEN JNCSAI. ISSN 0021-9533.

Othmer:1993:OWM

[361] H. G. Othmer and Y. Tang. Oscillations and waves in a model of ip-controlled calcium dynamics. In Othmer et al. [542], pages 277–313. ISBN 0-306-44661-8. LCCN QH491 .N38 1993.

Othmer:1993:TAB

[362] H. G. Othmer. Techniques and applications of bifurcation theory. Lecture Notes, University of Utah, 1993.

Pate:1993:DMS

[363] Edward Pate, Howard D. White, and Roger Cooke. Determination of the myosin step size from mechanical and kinetic data. Proceedings of the National Academy of Sciences of the United States of America, 90(6):2451–2455, March 15, 1993. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Pate:1993:UDN

[364] E. Pate, K. Franks-Skiba, H. D. White, and R. Cooke. The use of differing nucleotides to investigate cross-bridge kinetics. Journal of biological chemistry, 268(14):10046–10053, May 15, 1993. CODEN JBCHA3. ISSN 0021-9258.

Peskin:1993:CBD

[365] C. S. Peskin and D. M. McQueen. Computational biofluid dynamics. Contemporary mathematics, 141:161–186, 1993. ISSN 0271-4132 (print), 1098-3627 (electronic).

Ross:1993:SSM

[366] A. H. Ross, W. S. C. Gurney, M. R.

Heath, S. J. Hay, and E. W. Henderson. A strategic simulation model of a fjord ecosystem. Limnology and Oceanography, 38(1):128–153, January 1993. CO-DEN LIOCAH. ISSN 0024-3590.

Sharp:1993:DCA

[367] Andrew A. Sharp, Michael B. O’Neil, L. F. Abbott, and Eve Marder. The dynamic clamp: artificial conductances in biological neurons. Trends in Neurosciences (Regular ed.), 16(10):389–394, October 1993. CODEN TNSCDR. ISSN 0378-5912.

Sharp:1993:DCC

[368] A. A. Sharp, M. B. O’Neil, L. F. Abbott, and E. Marder. Dynamic clamp: Computer-generated conductances in real neurons. Journal of Neurophysiology, 69(3):992–995, March 1993. CODEN JONEA4. ISSN 0022-3077.

Starmer:1993:VEM

[369] C. F. Starmer, V. N. Biktahev, D. N. Romashko, M. R. Stepanov, O. N. Makarova, and V. I. Krinsky. Vulnerability in an excitable medium: analytical and numerical study of initiating unidirectional propagation. Biophysical Journal, 65(5):1775–1787, November 1993. CODEN BIOJAU. ISSN 0006-3495.

Templeton:1993:HGC

[370] A. R. Templeton. The “Eve” hypothesis: a genetic critique and reanalysis. American Anthropologist, 95:51–72, 1993. CO-DEN AMATA7. ISSN 0002-7294.

Trayanova:1993:BMP

[371] N. A. Trayanova and T. C. Pilkington. A bidomain model with periodic intracellular junctions. one-dimensional analysis. IEEE Transactions on Biomedical Engineering, 40:424–433, 1993. CODEN IEBEAX. ISSN 0018-9294.

Tuljapurkar:1993:EUV

[372] S. Tuljapurkar and C. Istock. Environmental uncertainty and variable diapause. Theoretical Population Biology, 43:251–280, 1993. CODEN TLPBAQ. ISSN 0040-5809.

White:1993:KBH

[373] H. D. White, B. Belknap, and W. Jiang.

Kinetics of binding and hydrolysis of a series of nucleoside triphosphates by actomyosin-S1: relationship between solution rate constants and properties of muscle fibers. Journal of biological chemistry, 268(14):10039–10045, May 15, 1993. CODEN JBCHA3. ISSN 0021-9258.

Zhang:1993:CSM

[374] X. Zhang, J. A. Ashton-Miller, and C. S. Stohler. A closed-loop system for maintaining constant experimental muscle pain in man. IEEE Transactions on Biomedical Engineering, 40(4):344–352, April 1993. CODEN IEBEAX. ISSN 0018-9294.

Zhou:1993:EMV

[375] X. Zhou, P. Daubert, P. D. Wolf, W. M.

Smith, and R. E. Ideker. Epicardial mapping of ventricular defibrillation with monophasic and biphasic shocks in dogs. Circulation Research, 72:145–160, 1993. CODEN CIRUAL. ISSN 0009-7330.

Zipser:1993:SNM

[376] D. Zipser, B. Kehoe, G. Littlewort, and J. Fuster. A spiking network model of short-term memory. Journal of Neuroscience, 13(8):3406–3420, August 1993. CODEN JNRSDS. ISSN 0270-6474.

Alberts:1994:MBC

[377] B. Alberts, D. Bray, J. Lewis, M. Raff, K. Roberts, and J. D. Watson. Molecular Biology of the Cell. Garland Publishing, Inc., New York, NY, USA, third edition, 1994. ISBN 0-8153-1619-4 (hardcover), 0-8153-1620-8 (paperback). xliii + 1294 + 67 pp. LCCN QH581.2 .M64 1994.

Altes:1994:EPP

[378] H. Korthals Altes. Effects of PDV on the population dynamics of seals. Master’s thesis, Leiden University, Leiden, The Netherlands, 1994.

Amon:1994:CCC

[379] A. Amon, S. Irniger, and K. Nasmyth. Closing the cell cycle circle in yeast: G2 cyclin proteolysis initiated at mitosis persists until the activation of G1 cyclins in the next cycle. Cell, 77:1037– 1050, 1994. CODEN CELLB5. ISSN 0092-8674.

Bezprozvanny:1994:IIC

[380] I. Bezprozvanny and B. Ehrlich. Inositol (1,4,5)-trisphosphate (InsP_{3})-gated Ca channels from cerebellum: Conduction properties for divalent cations and regulation by intraluminal calcium. Journal of General Physiology, 104:821–856, 1994. CODEN JGPLAD. ISSN 0022-1295.

Chesson:1994:MCV

[381] P. L. Chesson. Multispecies competition in varying environments. Theoretical Population Biology, 45:227–276, 1994. CODEN TLPBAQ. ISSN 0040-5809.

Cooke:1994:MRM

[382] R. Cooke, H. D. White, and E. Pate. A model of the release of myosin heads from actin in rapidly contracting muscle. Biophysical Journal, 66(3):778–788, March 1994. CODEN BIOJAU. ISSN 0006-3495.

Ellner:1994:ROG

[383] S. Ellner and N. G. Hairston, Jr. Role of overlapping generations in maintaining genetic variation in a fluctuating environment. American Naturalist, 143:403– 417, 1994. CODEN AMNTA4. ISSN 0003-0147.

Finer:1994:SMM

[384] J. T. Finer, R. M. Simmons, and J. A. Spudich. Single myosin molecule mechanics: PicoNewton forces and nanometre steps. Nature, 368(6467): 113–119, March 10, 1994. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Freeman:1994:TEP

[385] S. A. Freeman, M. A. Wang, and J. C.

Weaver. Theory of electroporation of planar bilayer membranes: Predictions of the aqueous area, change in capacitance, and pore-pore separation. Biophysical Journal, 67:42–56, 1994. CO-DEN BIOJAU. ISSN 0006-3495.

Glass:1994:PFL

[386] L. Glass and J. Sun. Periodic forcing of a limit cycle oscillator: Fixed points, arnold tongues, and the global organization of bifurcations. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 50: 5077–5084, 1994. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic).

Griffiths:1994:AIP

[387] R. C. Griffiths and S. Tavaré. Ancestral inference in population genetics. Statistical Science, 9:307–319, 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Griffiths:1994:STN

[388] R. C. Griffiths and S. Tavaré. Sampling theory for neutral alleles in a varying environment. Philosophical Transactions of the Royal Society of London Series B, Biological sciences, 344:403–410, 1994. CODEN PTRBAE. ISSN 0962-8436.

Kowtha:1994:IMN

[389] V. C. Kowtha, A. Kunysz, J. R. Clay, L. Glass, and A. Shrier. Ionic mechanisms and nonlinear dynamics of embryonic chick heart cell aggregates. Progress in Biophysics and Molecular Biology, 61: 255–281, 1994. CODEN PBIMAC. ISSN 0079-6107.

Krassowska:1994:EBC

[390] W. Krassowska and J. C. Neu. Effective boundary conditions for syncytial tissues. IEEE Transactions on Biomedical Engineering, 41:143–150, 1994. CO-DEN IEBEAX. ISSN 0018-9294.

Krassowska:1994:RSC

[391] W. Krassowska and J. C. Neu. Response of a single cell to an external electric field. Biophysical Journal, 66:1768– 1776, 1994. CODEN BIOJAU. ISSN 0006-3495.

Middlebrooks:1994:PCS

[392] A. E. Middlebrooks, L. Xu, and D. M. Green. A panoramic code for sound location by cortical neurons. Science, 264 (5160):842–844, May 6, 1994. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Moreno:1994:RPT

[393] S. Moreno and P. Nurse. Regulation of progression through the G1 phase of the cell cycle by the rum1^{+} gene. Nature, 367(6460):236–242, January 20, 1994. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Peskin:1994:MED

[394] C. S. Peskin and D. M. McQueen.

Mechanical equilibrium determines the fractal fiber architecture of the aortic heart valve leaflets. American Journal of Physiology, 266(1):H319–H328, January 1994. CODEN AJPHAP. ISSN 0002-9513 (print), 2163-5773 (electronic).

Ross:1994:CSE

[395] A. H. Ross, W. S. C. Gurney, and M. R.

Heath. A comparative study of the ecosystem dynamics of four fjords. Limnology and Oceanography, 39(2):318– 343, March 1994. CODEN LIOCAH. ISSN 0024-3590.

Schiff:1994:CCB

[396] S. J. Schiff, K. Jerger, D. H. Duong, T. Chang, M. L. Spano, and W. L. Ditto. Controlling chaos in the brain. Nature, 370(6491):615–620, August 25, 1994. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Siegfried:1994:BCE

[397] J. Siegfried. Bilateral chronic electrostimulation of ventroposteriolateral pallidum: a new therapeutic approach for alleviating all parkinsonian symptoms. Neurosurgery, 35(6):1126–1129, 1994. ISSN 0148-396X.

Tang:1994:GMA

[398] Yuanhua Tang and Hans G. Othmer. A G-protein-based model of adaptation in Dictyostelium discoideum. Mathematical Biosciences, 120:25–76, 1994. CODEN MABIAR. ISSN 0025-5564.

Tang:1994:MCD

[399] Yuanhua Tang and Hans G. Othmer.

A model of calcium dynamics in cardiac myocytes based on the kinetics of ryanodine-sensitive calcium channels. Biophysical Journal, 67:2223–2235, 1994. CODEN BIOJAU. ISSN 0006-3495.

Ayala:1995:AAT

[400] F. J. Ayala. Association affairs: The myth of Eve: Molecular biology and human origins. Science, 270:1930–1936, 1995. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Behbehani:1995:ACA

[401] K. Behbehani, F-C. Yen, J. R. Burk, E. A. Lucas, and J. R. Axe. Automatic control of airway pressure for treatment of obstructive sleep apnea. IEEE Transactions on Biomedical Engineering, 42(10):1007–1016, 1995. CO-DEN IEBEAX. ISSN 0018-9294.

Bulmer:1995:TEE

[402] Michael Bulmer. Theoretical Evolutionary Ecology. Sinauer Associates, Inc., Sunderland, MA, 1995. ISBN 0-87893-078-7 (paper), 0-87893-079-5 (cloth). xi + 352 pp. LCCN QH366.2 .B85 1994.

Burger:1995:EEC

[403] R. Burger and M. Lynch. Evolution and extinction in a changing environment: a quantitative-genetic analysis. Evolution, 49:151–163, 1995. CODEN EVOLAO. ISSN 0014-3820.

Burt:1995:PEF

[404] A. Burt. Perspective: The evolution of fitness. Evolution, 49:1–8, 1995. CO-DEN EVOLAO. ISSN 0014-3820.

Campbell:1995:CDM

[405] S. A. Campbell, J. Bélair, T. Ohira, and J. Milton. Complex dynamics and multistability in a damped harmonic oscillator with delayed negative feedback. Chaos (Woodbury, NY), 5:640– 645, 1995. CODEN CHAOEH. ISSN 1054-1500.

Correa-Bordes:1995:POS

[406] J. Correa-Bordes and P. Nurse. p25rum1 orders S-phase and mitosis by acting as an inhibitor of the p34cdc2 mitotic kinase. Cell, 83:1001–1009, 1995. CODEN CELLB5. ISSN 0092-8674.

DiBenedetto:1995:PDE

[407] Emmanuele DiBenedetto. Partial Differential Equations. Birkhäuser Boston Inc., Cambridge, MA, USA, 1995. ISBN 0-8176-3708-7 (Boston), 3-7643-3708-7 (Basel). xiv + 416 pp. LCCN QA377 .D624 1995.

Diekmann:1995:FIP

[408] O. Diekmann, M. C. M. de Jong, A. A.

de Koeijer, and P. Reijnders. The force of infection in populations of varying size: A modelling problem. Journal of Biological Systems, 3:519–529, 1995. CODEN JBSYE2. ISSN 0218-3390.

Donnelly:1995:CGS

[409] P. Donnelly and S. Tavaré. Coalescents and genealogical structure under neutrality. Annual Review of Genetics, 29: 401–421, 1995. CODEN ARVGB7. ISSN 0066-4197.

Dorit:1995:APA

[410] R. L. Dorit, H. Akashi, and W. Gilbert.

Absence of polymorphism at the ZFY locus on the human Y chromosome. Science, 268:1183–1185, 1995. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Easterling:1995:ODS

[411] M. E. Easterling. Optimal dormancy strategies in structured population. Master’s thesis, North Carolina State University, Raleigh, NC, 1995.

Ferster:1995:CNC

[412] D. Ferster and N. Spruston. Cracking the neuronal code. Science, 270(5237): 756–757, November 3, 1995. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Fromherz:1995:SJC

[413] Peter Fromherz and Alfred Stett. Silicon-neuron junction: Capacitive stimulation of an individual neuron on a silicon chip. Physical Review Letters, 75 (8):1670–1673, 1995. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Furutani:1995:BPC

[414] E. Furutani, M. Araki, T. Sakamoto, and S. Maetani. Blood pressure control during surgical operations. IEEE Transactions on Biomedical Engineering, 42 (10):999–1006, 1995. CODEN IEBEAX. ISSN 0018-9294.

Glass:1995:RAR

[415] L. Glass and M. E. Josephson. Resetting and annihilation of reentrant abnormally rapid heartbeat. Physical Review Letters, 75(10):2059–2063, 1995. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Gomulkiewicz:1995:WDE

[416] R. Gomulkiewicz and R. D. Holt. When does evolution by natural selection prevent extinction? Evolution, 49:201–207, 1995. CODEN EVOLAO. ISSN 0014-3820.

Grenfell:1995:EID

[417] B. T. Grenfell and A. Dobson, editors.

Ecology of Infectious Diseases in Natural Populations. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-46502-8. xii + 521 pp. LCCN QR100 .E32 1995.

Griffiths:1995:UGT

[418] R. C. Griffiths and S. Tavaré. Unrooted genealogical tree probabilities in the infinitely-many-sites model. Mathematical Biosciences, 127:77–98, 1995. CODEN MABIAR. ISSN 0025-5564.

Grover:1995:CHE

[419] J. P. Grover. Competition, herbivory, and enrichment-nutrient-based models for edible and inedible plants. American Naturalist, 145(5):746–774, May 1995. CODEN AMNTA4. ISSN 0003-0147.

Haccou:1995:OMS

[420] P. Haccou and Y. Iwasa. Optimal mixed strategies in stochastic environments. Theoretical Population Biology, 47:212– 243, 1995. CODEN TLPBAQ. ISSN 0040-5809.

Hairston:1995:ASD

[421] N. G. Hairston, Jr., R. A. van Brunt, C. M. Kearns, and D. R. Engstrom. Age and survivorship of diapausing eggs in a sediment egg bank. Ecology, 76:1706– 1711, 1995. CODEN ECOLAR. ISSN 0012-9658.

Hammer:1995:RCA

[422] M. F. Hammer. A recent common ancestry for human Y chromosomes. Nature, 378(6555):376–378, 1995. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Herwaarden:1995:SEM

[423] O. A. van Herwaarden and J. Grasman. Stochastic epidemics: Major outbreaks and the duration of the epidemic. Journal of Mathematical Biology, 33:581– 601, 1995. CODEN JMBLAJ. ISSN 0303-6812.

Higuchi:1995:CTF

[424] H. Higuchi, T. Yanagida, and Y. E. Goldman. Compliance of thin filaments in skinned fibers of rabbit skeletal muscle. Biophysical Journal, 69(3):1000– 1010, 1995. CODEN BIOJAU. ISSN 0006-3495.

Hillesdon:1995:DCG

[425] A. J. Hillesdon, T. J. Pedley, and J. O. Kessler. The development of concentration gradients in a suspension of chemotactic bacteria. Bulletin of Mathematical Biology, 57(2):299–344, 1995. CODEN BMTBAP. ISSN 0092-8240.

Iacono:1995:SGP

[426] R. P. Iacono. Stimulation of the globus pallidus in parkinson’s disease. British Journal of Neurosurgery, 9:505– 510, 1995. CODEN BJNEEL. ISSN 0268-8697.

Jobling:1995:FSY

[427] M. A. Jobling and C. Tyler-Smith. Fathers and sons: the Y chromosome and human evolution. Trends in Genetics, 11:449–456, 1995. CODEN TRGEE2. ISSN 0168-9525.

Kawecki:1995:DSP

[428] T. J. Kawecki. Demography of sourcesink populations and the evolution of ecological niches. Evolutionary Ecology, 7:155–174, 1995. CODEN EVECEJ. ISSN 0269-7653.

Kostov:1995:MLC

[429] A. Kostov, B. J. Andrews, D. B.

Popovic, R. B. Stein, and W. W. Armstrong. Machine learning in control of functional electrical stimulation systems for locomotion. IEEE Transactions on Biomedical Engineering, 42(6):541–551, June 1995. CODEN IEBEAX. ISSN 0018-9294.

Krassowska:1995:EET

[430] W. Krassowska. Effects of electroporation on transmembrane potential induced by defibrillation shocks. Pacing and Clinical Electrophysiology: PACE, 18:1644–1660, 1995. ISSN 0147-8389.

Krassowska:1995:MME

[431] W. Krassowska. Macroscopic model of electroporating membrane. In IEEE [544], pages 253–254.

Kunysz:1995:OSS

[432] A. Kunysz, L. Glass, and A. Shrier.

Overdrive suppression of spontaneously beating chick heart cell aggregates: Experiment and theory. American Journal of Physiology, 269:H1153–H1164, 1995. CODEN AJPHAP. ISSN 0002-9513 (print), 2163-5773 (electronic). (Heart Circ. Physiol. 38).

Kuznetsov:1995:EAB

[433] Y. A. Kuznetsov. Elements of Applied Bifurcation Theory. Springer-Verlag, 1995.

Lodish:1995:MCB

[434] Harvey F. Lodish, D. Baltimore, A. Berk, S. L. Zipursky, P. Matsudaira, and James E. Darnell. Molecular Cell Biology. W. H. Freeman, New York, NY, USA, third edition, 1995. ISBN 0-7167-2380-8. various pp. LCCN QH581.2 .D37 1995.

Lucas:1995:RJ

[435] G. Lucas. Return of the Jedi. Warner Books, New York, NY, USA, 1995.

Mainen:1995:RST

[436] Z. F. Mainen and T. J. Sejnowski. Reliablity of spike timing in neocortical neurons. Science, 268(5216):1503– 1506, June 9, 1995. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Maniglia:1995:MIE

[437] A. J. Maniglia (ed.). Middle and inner ear electronic implantable devices for partial hearing loss. The Otolaryngologic clinics of North America, 28:1–223, 1995. CODEN OCNABW. ISSN 0030-6665.

McNamara:1995:DOF

[438] J. M. McNamara, J. N. Webb, and E. J. Collins. Dynamic optimization in fluctuating environments. Proceedings of the Royal Society of London. Series B. Biological sciences, 261:279–284, 1995. CO-DEN PRLBA4. ISSN 0962-8452 (print), 1471-2954 (electronic).

McNamara:1995:IFD

[439] J. M. McNamara. Implicit frequency dependence and kin selection in fluctuating environments. Evolutionary Ecology, 9: 185–203, 1995. CODEN EVECEJ. ISSN 0269-7653.

Mensour:1995:CCS

[440] B. Mensour and A. Longtin. Controlling chaos to store information in delaydifferential equations. Physics Letters A, 205(1):18–24, 1995. CODEN PYLAAG. ISSN 0375-9601 (print), 1873-2429 (electronic).

Milton:1995:DFD

[441] J. Milton, S. A. Campbell, and J. Bélair.

Dynamic feedback and the design of closed-loop drug delivery systems. Journal of Biological Systems, 3:711–718, 1995. CODEN JBSYE2.

Mitchell-Olds:1995:MBQ

[442] T. Mitchell-Olds. The molecular basis of quantitative genetic variation in natural populations. Trends in Ecology and Evolution, 10:324–328, 1995. CODEN TREEEQ. ISSN 0169-5347.

Mollison:1995:EMT

[443] D. Mollison, editor. Epidemic Models: Their Structure and Relation to Data. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-47536-8. xvii + 424 pp. LCCN RA652.2.M3 E64 1995.

Molloy:1995:MFP

[444] J. E. Molloy, J. Kendrick-Jones, J. E.

Burns, R. T. Tregear, and D. C. S. White. Movement and force produced by a single myosin head. Nature, 378 (6553):209–212, November 9, 1995. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Murray:1995:CUD

[445] A. W. Murray. Cyclin ubiquitination:

the destructive end of mitosis. Cell, 81:

149–152, 1995. CODEN CELLB5. ISSN 0092-8674.

Novak:1995:QAM

[446] B. Novak and J. J. Tyson. Quantitative analysis of a molecular model of mitotic control in fission yeast. Journal of Theoretical Biology, 173:283–305, 1995. CO-DEN JTBIAP. ISSN 0022-5193.

Peskin:1995:GMC

[447] C. S. Peskin and D. M. McQueen. A general method for the computer simulation of biological systems interacting with fluids. In Ellington and Pedley [546], pages 265–276. ISBN 0-948601-50-7. LCCN QH301 .S74 no.49.

Rudy:1995:RIT

[448] Y. Rudy. Reentry: Insights from theoretical simulations in a fixed pathway. Journal of Cardiovascular Electrophysiology, 6(4):294–312, 1995. CODEN JCELE2. ISSN 1045-3873.

Sasaki:1995:ESP

[449] A. Sasaki and S. Ellner. The evolutionarily stable phenotype distribution in a random environment. Evolution, 49:336–350, 1995. CODEN EVOLAO. ISSN 0014-3820.

Sellers:1995:PPM

[450] James R. Sellers and Holly V. Goodson. Protein Profile. Motor Proteins 2: Myosin, volume 2(12) of Protein profile. Academic Press, New York, NY, 1995. ISSN 1070-3667. 1423 pp.

Smith:1995:TCD

[451] H. L. Smith and P. Waltman. The Theory of the Chemostat: Dynamics of Microbial Competition, volume 13 of Cambridge Studies in Mathematical Biology. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-47027-7. xvi + 313 pp. LCCN QR84.5 .S65 1995.

Sneyd:1995:MCO

[452] J. Sneyd, J. Keizer, and M. J. Sanderson. Mechanisms of calcium oscillations and waves: A quantitative analysis. The FASEB journal: official publication of the Federation of American Societies for Experimental Biology, 9:1463– 1472, 1995. CODEN FAJOEC. ISSN 0892-6638.

Sobie:1995:BSI

[453] E. A. Sobie, M. G. Fishler, and L. Tung.

Biphasic shocks induce more uniform repolarization than monophasic shocks in a cardiac cellular field excitation model. In IEEE [544], pages 257–258.

Tang:1995:FEF

[454] Yuanhua Tang and Hans G. Othmer.

Frequency encoding in forced excitable systems with applications to calcium oscillations. Proceedings of the National Academy of Sciences of the United States of America, 92:7869–7873, 1995. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Trayanova:1995:VEE

[455] N. Trayanova and W. Krassowska. Virtual electrode effects in a bidomain model with electroporating membrane. In IEEE [544], pages 255–256.

Tyson:1995:CCC

[456] J. J. Tyson, B. Novak, K. C. Chen, and J. Val. Checkpoints in the cell cycle from a modeler’s perspective. Progress in Cell Cycle Research, 1:1–8, 1995. ISSN 1087-

Wallace:1995:MDV

[457] D. C. Wallace. Mitochondrial DNA variation in human evolution, degenerative disease, and aging. American Journal of Human Genetics, 57:201–223, 1995. CO-DEN AJHGAG. ISSN 0002-9297.

Whitfield:1995:SVH

[458] L. S. Whitfield, J. E. Sulston, and P. N. Goodfellow. Sequence variation of the human Y chromosome. Nature, 378 (6555):379–380, 1995. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Widdows:1995:SGC

[459] J. Widdows, P. Donkin, M. D. Brinsley, S. V. Evans, P. N. Salkeld, A. Franklin, R. J. Law, and M. J. Waldock. Scope for growth and contaminant levels in north sea mussels Mytilus edulis. Marine Ecology Progress Series, 127(1–3):131–148, 1995. CODEN MESEDT. ISSN 0171-8630.

Wills:1995:WDE

[460] C. Wills. When did Eve live? an evolutionary detective story. Evolution, 49:593–607, 1995. CODEN EVOLAO. ISSN 0014-3820.

Windish:1995:OMM

[461] H. Windish, H. Ahammer, P. Schaffer, W. Müller, and D. Platzer. Optical multisite monitoring of cell excitation phenomena in isolated cardiomyocyte. Pflugers Archiv: European Journal of Physiology, 430(4):508–518, 1995. CO-DEN PFLABK. ISSN 0031-6768.

Zhou:1995:OTP

[462] X. Zhou, R. E. Ideker, T. F. Blitchington, W. M. Smith, and S. B. Knisley.

Optical transmembrane potential measurements during defibrillation-strength shocks in perfused rabbit hearts. Circulation Research, 77(3):593–602, 1995. CODEN CIRUAL. ISSN 0009-7330.

Zhou:1995:RTP

[463] X. Zhou, D. L. Rollins, W. M. Smith, and R. E. Ideker. Responses of the transmembrane potential of myocardial cells during a shock. Journal of Cardiovascular Electrophysiology, 6:252–263, 1995. CODEN JCELE2. ISSN 1045-3873.

Ball:1996:ETL

[464] F. Ball, D. Mollison, and G. Scalia-Tomba. Epidemics with two levels of mixing. Submitted., 1996.

Bayer:1996:OTB

[465] W. Bayer and U. an der Heiden. Oscillation types and bifurcations of a nonlinear second order differential-difference equation. Journal of Dynamics and Differential Equations, 1996. CODEN JD-DEEH. ISSN 1040-7294. In press.

Bees:1996:MBP

[466] M. A. Bees and N. A. Hill. Measurements of biconvection pattern wavelengths in suspension of Chlamydomonas. In preparation, 1996.

Bees:1996:WBP

[467] M. A. Bees and N. A. Hill. Wavelengths of biconvection pattern wavelengths in suspension of Chlamydomonas. In preparation, 1996.

Dallon:1996:DCM

[468] John. C. Dallon and Hans G. Othmer.

A discrete cell model with adaptive signalling for aggregation of Dictyostelium discoideum. To appear in Phil. Trans. Roy. Soc. Lon., 1996.

DeBruin:1996:ACI

[469] K. A. DeBruin and W. Krassowska. Electroporation as a mechanism of the saturation of transmembrane potential induced by large electric fields. In 18th Annual Conference of the IEEE Engineering in Medicine and Biology Society. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. Accepted.

Diekmann:1996:FSE

[470] O. Diekmann, A. A. de Koeijer, and J. A. J. Metz. On the final size of epidemics within herds. Canadian Applied Mathematics Quarterly, 4(1):21– 30, 1996. ISSN 1073-1849.

Donnelly:1996:TSA

[471] P. Donnelly, S. Tavaré, D. J. Balding, and R. C. Griffiths. On the time since Adam. Science, 272:1357–1359, 1996. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Ellner:1996:EFM

[472] S. Ellner. Environmental fluctuations and the maintenance of genetic diversity in age or stage-structured populations. Bulletin of Mathematical Biology, 58:103–128, 1996. CODEN BMTBAP. ISSN 0092-8240.

Ellner:1996:PGP

[473] S. Ellner and A. Sasaki. Patterns of genetic polymorphism maintained by fluctuating selection with overlapping generations. Theoretical Population Biology, 50:31–65, 1996. CODEN TLPBAQ. ISSN 0040-5809.

Foss:1996:MDR

[474] Jennifer Foss, Andre Longtin, Boualem Mensour, and John Milton. Multistability and delayed recurrent loops. Physical Review Letters, 76(4):708–711, 1996. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Foss:1996:NMD

[475] J. Foss, F. Moss, and J. Milton.

Noise, multistability and delayed recurrent loops. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 1996. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). Submitted.

Gurney:1996:EDS

[476] W. S. C. Gurney, D. A. J. Middleton, R. M. Nisbet, E. McCauley, W. W. Murdoch, and A. M. de Roos. The equilibrium demography of structured populations. Theoretical Population Biology, 1996. CODEN TLPBAQ. ISSN 0040-5809. In press.

Hairston:1996:OGS

[477] N. G. Hairston, Jr., S. Ellner, and C. M.

Kearns. Overlapping generations: the storage effect and the maintenance of biotic diversity. In Rhodes, Jr. et al. [549], pages 109–145. ISBN 0-226-71057-2 (cloth), 0-226-71058-0 (paper). LCCN QH352 .P628 1996.

Hanski:1996:MVP

[478] I. Hanski, A. Moilanen, and M. Gyllenberg. Minimum viable population size. American Naturalist, 147:527–541, 1996. CODEN AMNTA4. ISSN 0003-0147.

Herwaarden:1996:SET

[479] O. A. van Herwaarden. Stochastic epidemics: The probability of extinction of an infectious disease at the end of a major outbreak. Journal of Mathematical Biology, 1996. CODEN JMBLAJ. ISSN 0303-6812. In press.

Hill:1996:BRWa

[480] N. A. Hill and D.-P. Häder. A biased random walk model for the trajectories of swimming micro-organisms. Journal of Theoretical Biology, 1996. CODEN JTBIAP. ISSN 0022-5193. Submitted.

Hill:1996:BRWb

[481] N. A. Hill and D.-P. Hader. A biased random walk model for the trajectories of swimming micro-organisms. Submitted to Philosophical Transactions of the Royal Society of London B., 1996.

Hillesdon:1996:PFS

[482] A. J. Hillesdon and T. J. Pedley. Pattern formation in suspensions of oxytactic bacteria: liner theory. Journal of Fluid Mechanics, 1996. CODEN JFLSA7. ISSN 0022-1120. To appear.

Holt:1996:CPD

[483] R. D. Holt and M. A. McPeek. Chaotic population dynamics favors dispersal. American Naturalist, 1996. CODEN AMNTA4. ISSN 0003-0147. In press.

Holt:1996:DCE

[484] R. D. Holt. Demographic constraints in evolution: Towards unifying the evolutionary theories on senescence and niche conservatism. Evolutionary Ecology, 10: 1–11, 1996. CODEN EVECEJ. ISSN 0269-7653.

Holt:1996:HDI

[485] R. D. Holt and R. Gomulkiewicz. How does immigration influence local adaptation? A re-examination of a familiar paradigm. American Naturalist, 1996. CODEN AMNATH. ISSN 0269-7699. In press.

Keener:1996:DAD

[486] J. P. Keener. Direct activation and defibrillation of cardiac tissue. Journal of Theoretical Biology, 178(313–324), 1996. CODEN JTBIAP. ISSN 0022-5193.

Koeijer:1996:MSP

[487] A. A. de Koeijer, O. Diekmann, and P. Reijnders. Modelling the spread of Phocine Distemper Virus (PDV) among harbor seals. Submitted., 1996.

Koutalos:1996:RSV

[488] Y. Koutalos and K. Y. Yau. Regulation of sensitivity in vertebrate rod photoreceptors by calcium. Trends in Neurosciences (Regular ed.), 19(2):73– 81, 1996. CODEN TNSCDR. ISSN 0166-2236, 0378-5912.

Krassowska:1996:BMPa

[489] W. Krassowska. The bidomain model.

part I: Derivation. IEEE Transactions on Biomedical Engineering, 1996. CO-DEN IEBEAX. ISSN 0018-9294. Submitted.

Krassowska:1996:BMPb

[490] W. Krassowska. The bidomain model.

part II: Limitations. IEEE Transactions on Biomedical Engineering, 1996. CO-DEN IEBEAX. ISSN 0018-9294. Submitted.

Krassowska:1996:RSI

[491] W. Krassowska and M. S. Kumar. The role of spatial interactions in creating the dispersion of transmembrane potential by premature electric shocks. Annals of biomedical engineering, 1996. CO-DEN ABMECF. ISSN 0090-6964. Submitted.

Kunysz:1996:BBD

[492] A. Kunysz, A. Shrier, and L. Glass. Bursting behavior during fixed delay stimulation of spontaneously beating chick heart cell aggregates. American Journal of Physiology, 1996. CODEN AJPHAP. ISSN 0002-9513 (print), 2163-5773 (electronic). Submitted.

Lebert:1996:HET

[493] M. Lebert and D.-P. Häder. How Euglena tells up from down. Nature, 379 (6566):590, 1996. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

McCauley:1996:SPM

[494] E. McCauley, R. M. Nisbet, A. M. de Roos, W. W. Murdoch, and W. S. C. Gurney. Structured population models of herbivorous zooplankton. Ecology, 1996. CODEN ECOLAR. ISSN 0012-9658. In press.

McNamara:1996:OLH

[495] J. M. McNamara. Optimal life histories for structured populations in fluctuating environments. Theoretical Population Biology, 1996. CODEN TLPBAQ. ISSN 0040-5809. In press.

Milton:1996:MPL

[496] J. Milton, W. Bayer, and U. an der Heiden. Modeling the pupil light reflex with differential delay equations. Zeitschrift für Angewandte Mathematik und Mechanik, 1996. CODEN ZAM-MAX. ISSN 0044-2267 (print), 1521-4001 (electronic). In press.

Murdoch:1996:FPE

[497] W. W. Murdoch and R. M. Nisbet. Frontiers of population ecology. In Floyd and Sheppard [548].

Murdoch:1996:PAD

[498] W. W. Murdoch, R. M. Nisbet, E. Mc-Cauley, A. M. de Roos, and W. S. C. Gurney. Plankton abundance and dynamics across nutrient levels: Tests of hypotheses. Submitted to Ecology., 1996.

Nisbet:1996:MRI

[499] R. M. Nisbet, E. B. Muller, A. J. Brooks, and P. Hosseini. Models relating individual and population response to contaminants. Submitted to Environmental Modeling and Assessment., 1996.

Nomura:1996:ETR

[500] T. Nomura and L. Glass. Entrainment and termination of reentrant wave propagation in a periodically stimulated ring of excitable media. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 53: 6353–6360, 1996. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic).

Pumir:1996:HDE

[501] A. Pumir and V. I. Krinsky. How does an electric field defibrillate cardiac muscle? Physica D, 91:205–219, 1996. CODEN PDNPDT. ISSN 0167-2789 (print), 1872-8022 (electronic).

Reijnders:1996:RSC

[502] P. J. H. Reijnders, S. M. J. M. Brasseur, and E. H. Ries. The release of seals from captive breeding and rehabilitation programmes: a useful conservation management tool? In Aubin et al. [547], pages 54–65. ISBN ???? LCCN ????

Richardson:1996:DBS

[503] D. E. Richardson and J. R. Goreck. Deep brain stimulation for pain relief. In Wilkins and Rengachary [550], pages 4021–4027. ISBN 0-07-079991-1 (set: hard cover), 0-07-070313-2 (vol. 1: hard cover) 0-07-070314-0 (vol. 2: hard cover) 0-07-070315-9 (vol. 3: hard cover). LCCN WL368 N52 1996.

Roos:1996:RIL

[504] A. M. de Roos, E. McCauley, R. M. Nisbet, W. S. C. Gurney, and W. W. Murdoch. Relating individual life-history and population dynamics in Daphnia pulex: Is the population more than a collection of individuals?, 1996. In preparation.

Sasaki:1996:QGV

[505] A. Sasaki and S. Ellner. Quantitative genetic variance and covariance maintained by fluctuating selection with overlapping generations. Evolution, 1996. CODEN EVOLAO. ISSN 0014-3820. Submitted.

Spiro:1996:MEA

[506] Peter Spiro, John S. Parkinson, and H. G. Othmer. A model of excitation and adaptation in bacterial chemotaxis. Submitted to PNAS, 1996.

Tang:1996:SAM

[507] Y. Tang, J. Stephenson, and H. G.

Othmer. Simplification and analysis of models of calcium dynamics based on ip-sensitive calcium channel dynamics. Biophysical Journal, 70:246–263, 1996. CODEN BIOJAU. ISSN 0006-3495.

Tavare:1996:ICT

[508] S. Tavaré, D. J. Balding, R. C. Griffiths, and P. Donnelly. Inferring coalescence times from DNA sequence data. Genetics, 1996. CODEN GENTAE. ISSN 0016-6731. Submitted.

Tuljapurkar:1996:SMM

[509] Shripad Tuljapurkar and Hal Caswell.

Structured-Population Models in Marine, Terrestrial, and Freshwater Systems. Chapman and Hall, Ltd., London, UK, 1996. ISBN 0-412-07261-0 (hardcover), 0-412-07271-8 (paperback). LCCN QH352 .S86 1996.

Tyson:1996:CKT

[510] J. J. Tyson, B. Novak, G. M. Odell, K. Chen, and C. D. Thron. Chemical kinetic theory as a tool for understanding the regulation of M-phase promoting factor in the cell cycle. Trends in biochemical sciences, 21:89–96, 1996. CO-DEN TBSCDB, TBSRDM. ISSN 0968-0004, 0167-7640, 0376-5067.

Vincent:1996:BSP

[511] R. V. Vincent and N. A. Hill. Bioconvection in a suspension of phototactic algae. Journal of Fluid Mechanics, 300:1– 29, 1996. CODEN JFLSA7. ISSN 0022-1120. To appear.

Zhou:1996:TPC

[512] X. Zhou, W. M. Smith, R. E. Ideker, and D. L. Rollins. Transmembrane potential changes caused by shocks in guinea pig papillary muscle. American Journal of Physiology, 1996. CODEN AJPHAP. ISSN 0002-9513 (print), 2163-5773 (electronic). to appear.

Hairston:1997:PVZ

[513] N. G. Hairston, Jr., C. M. Kearns, and S. Ellner. Phenotypic variation in a zooplankton egg bank. Ecology, 1997. CO-DEN ECOLAR. ISSN 0012-9658. In press.

Holt:1997:RES

[514] R. D. Holt. Rarity and evolution; some theoretical considerations. In Kunin and Gaston [551], pages 209–234. ISBN ???? LCCN ???? In press.

Leibovic:1969:IPN

[515] K. N. Leibovic, editor. Information processing in the nervous system; proceedings of a symposium held at the State University of New York at Buffalo, 21st–24th October, 1968. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969.

Green:1970:LED

[516] A. E. Green and J. E. Adkins. Large elastic deformations. Clarendon Press, Oxford, UK, 1970. ISBN 0-19-853334-9. xiv + 324 pp. LCCN QA935.G73 1970.

Rice:1972:SMN

[517] Stuart A. Rice, Karl F. Freed, and John C. Light, editors. Statistical mechanics; new concepts, new problems, new applications. University of Chicago Press, Chicago, IL, USA, 1972. ISBN 0-226-71115-3. LCCN QC175 .I5 1971. IU-PAP Conference on Statistical Mechanics (6th: 1971: Chicago, IL).

Fitzgibbon:1977:ND

[518] W. E. Fitzgibbon, III and H. F. Walker, editors. Nonlinear Diffusion, volume 14 of Research notes in mathematics. Pitman Publishing Ltd., London, UK, 1977. ISBN 0-273-01066-2. 232 pp. LCCN QA274.75 .N66.

Baan:1978:CSD

[519] Jan Baan, Abraham Noordergraaf, and Jeff Raines, editors. Cardiovascular System Dynamics. MIT Press, Cambridge, MA, USA, 1978. ISBN 0-262-18078-2. xiv + 618 pp. LCCN RC670.C37.

Segel:1980:MMM

[520] Lee A. Segel, editor. Mathematical Models in Molecular and Cellular Biology. Cambridge University Press, Cambridge, 1980. ISBN 0-521-22925-1. ix + 757 pp. LCCN QH506 .M38.

Reijnders:1981:MMW

[521] P. J. H. Reijnders and W. J. Wolff, editors. Marine Mammals of the Wadden Sea: final report of the section ‘marine mammals’ of the Wadden Sea Working Group. Balkema, Rotterdam, 1981. ISBN 90-6191-057-9. 64 pp. LCCN QL713.2 .M364 1982.

Strathern:1981:MBY

[522] Jeffrey N. Strathern, Elizabeth W.

Jones, and James R. Broach, editors. The Molecular Biology of the Yeast Saccharomyces. Cold Spring Harbor Laboratory, Cold Spring Harbor, NY, USA, 1981. ISBN 0-87969-149-2. x + 680 pp. LCCN QK623.S23 M643 1982.

Glover:1981:GTM

[523] S. W. Glover and D. A. Hopwood, editors. Genetics as a tool in microbiology. Cambridge University Press, Cambridge, UK, 1981. ISBN 0-521-23748-3. LCCN QR1 .S7 v.31.

Koch:1982:EPS

[524] G. Koch and F. Spizzichino, editors. Exchangeability in probability and statistics: proceedings of the International Conference on Exchangeability in Probability and Statistics, Rome, 6th–9th April, 1981, in honour of Professor Bruno de Finetti. North-Holland Publishing Co., Amsterdam, The Netherlands, 1982. ISBN 0-444-86403-2. LCCN QA273.A1 I55 1981.

Nurse:1984:MCC

[525] Paul Nurse and Eva Streiblova, editors. The Microbial Cell Cycle. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1984. ISBN 0-442-30778-0, 0-17-771103-5. 88 pp. LCCN QR73.7 .M53 1984.

Field:1985:OTW

[526] Richard J. Field and Maria Burger, editors. Oscillations and Traveling Waves in Chemical Systems. Wiley, 1985. ISBN 0-471-89384-6. LCCN QD502 .O83 1985. US$85.00.

Meinardus:1985:DEA

[527] G. Meinardus and G. Nurnberger, editors. Delay Equations, Approximation and Application: international symposium at the University of Mannheim, October 8–11, 1984, volume 74 of International series of numerical mathematics. Birkhäuser Verlag, Basel, Switzerland, 1985. ISBN 3-7643-1733-7. LCCN QA371 .D37 1985.

Press:1986:NRA

[528] William H. Press, Brian P. Flannery, Saul A. Teukolsky, and William T. Vetterling. Numerical Recipes: The Art of Scientific Computing. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-30811-9. 390–396; 449–453; and 646–647 (of xx + 818) pp. LCCN QA76.73.F25N84 1986.

Rensing:1987:TDH

[529] L. Rensing, U. an der Heiden, and M. C. Mackey, editors. Temporal Disorder in Human Oscillatory Systems, volume 36 of Springer series in synergetics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-17765-5. ix + 259 pp. LCCN QP84.6 .T44 1987.

Eisen:1988:SIC

[530] E. J. Eisen, M. M. Goodman, G. Namkoong, and B. S. Weir, editors. The Second International Conference on Quantitative Genetics. Sinauer Associates, Inc., Sunderland, MA, 1988. ISBN 0-87893-900-8, 0-87893-901-6 (paperback). xii + 724 pp.

Neumann:1989:EEC

[531] E. Neumann, A. E. Sowers, and C. A.

Jordan, editors. Electroporation and Electrofusion in Cell Biology. Plenum Press, New York, NY, USA; London, UK, 1989. ISBN 0-306-43043-6. xviii + 436 pp. LCCN QH585.5.E48 E44 1989.

Zipes:1990:CEC

[532] D. P. Zipes and J. Jalife, editors. Cardiac Electrophysiology: From Cell to Bedside. W. B. Saunders, Philadelphia, PA, USA, 1990. ISBN 0-7216-1408-6. xxiv + 1034 + 2 pp. LCCN WG330 C38 1990.

Borsellino:1990:ST

[533] A. Borsellino, L. Cervetto, and V. Torre, editors. Sensory Transduction. Plenum Press, New York, NY, USA; London, UK, 1990.

Goldstein:1991:DEA

[534] Jerome A. Goldstein, Franz Kappel, and Wwilhelm Schappacher, editors. Differential Equations with Applications in Biology, Physics, and Engineering. Marcel Dekker, Inc., New York, NY, 1991. ISBN 0-8247-8571-1. xiii + 332 pp. LCCN QA370 .D56 1991.

Guyton:1991:TMP

[535] Arthur C. Guyton. Textbook of Medical Physiology. W. B. Saunders Company, Philadelphia PA, 1991. ISBN 0-7216-3087-1. 97–109 (of xli + 1014) pp. LCCN QP34.5.G9 1991.

Holden:1991:NWP

[536] A. V. Holden, M. Markus, and Hans G. Othmer, editors. Nonlinear Wave Processes in Excitable Media, volume 244 of NATO ASI series. Series B, Physics. Plenum Press, New York, NY, USA; London, UK, 1991. ISBN 0-306-43800-3. LCCN QA927 .N27 1989.

DeAngelis:1992:IMA

[537] Donald L. DeAngelis and Louis J. Gross, editors. Individual-based models and approaches in ecology: populations, communities, and ecosystems. Routledge, Chapman and Hall, New York, NY, 1992. ISBN 0-412-03161-2 (hardcover), 0-412-03171-X (paperback). xix + 525 pp. LCCN QH352 .I53 1992. US$65.00 (hardcover), US$29.00 (paperback).

Takahata:1992:MME

[538] Naoyuki Takahata and Andrew G.

Clark, editors. Mechanisms of Molecular Evolution: Introduction to Molecular Paleopopulation Biology. Sinauer Associates, Inc., Sunderland, MA, 1992. ISBN 4-7622-6718-X, 0-87893-825-7. x + 250 pp. LCCN QH352 .M43 1993.

Chorin:1993:MIF

[539] Alexandre J. Chorin and Jerrold E.

Marsden. A Mathematical Introduction to Fluid Mechanics. Springer-Verlag, New York NY, 1993. ISBN 0-387-97918-2. ix + 169 pp. LCCN QA901.C53 1993.

Josephson:1993:TMM

[540] M. E. Josephson and H. J. J. Wellens, editors. Tachycardias: Mechanisms and Management, volume 6 of The Bakken Research Center series. Futura, Mount Kisco, NY, USA, 1993. ISBN 0-87993-549-9. xii + 564 pp. LCCN WG 330 T117-3 1993.

Kareiva:1993:BIG

[541] P. M. Kareiva, J. G. Kingsolver, and R. B. Huey, editors. Biotic Interactions and Global Climate Change. Sinauer Associates, Inc., Sunderland, MA, 1993. ISBN 0-87893-429-4, 0-87893-430-8 (paperback). xii + 559 pp. LCCN QH546 .B55 1993. US$55.00, US$35.00 (paperback).

Othmer:1993:ETA

[542] H. G. Othmer, P. K. Maini, and J. D. Murray, editors. Experimental and Theoretical Advances in Biological Pattern Formation, volume 259 of NATO ASI series. Series A, Life sciences. Plenum Press, New York, NY, USA; London, UK, 1993. ISBN 0-306-44661-8. ix + 388 pp. LCCN QH491 .N38 1993.

Yoshimura:1993:ASE

[543] J. Yoshimura and C. W. Clark, editors. Adaptation in Stochastic Environments, volume 98 of Lecture Notes in Biomathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-56681-3 (Berlin), 0-387-56681-3 (New York). vi + 193 pp. LCCN QH546 .A28 1993.

IEEE:1995:PAI

[544] IEEE, editor. Proceedings of the 17th Annual International Conference of the IEEE Engineering in Medicine and Biology Society (EMBS): Montreal, Canada, September 20–23, 1995. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995.

Bock:1995:CWG

[545] Gregory Bock and Kate Ackrill, editors. Calcium Waves, Gradients and Oscillations, volume 188 of Ciba Foundation symposium. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1995. ISBN 0-471-95234-6. LCCN QP535.C2 C267 1995. UK£47.50. Proceedings of a symposium held at Ciba Foundation, London, 26–28 April 1994.

Ellington:1995:BFD

[546] C. P. Ellington and T. J. Pedley, editors.

Biological Fluid Dynamics: Proceedings of a meeting held at the University of Leeds, UK, 4–8 July 1994, volume 49 of Symposia of the Society for Experimental Biology. The Company of Biologists Limited, Cambridge, UK, 1995. ISBN 0-948601-50-7. LCCN QH301 .S74 no.49.

Aubin:1996:RRR

[547] David J. St. Aubin, Joseph R. Geraci, and Valerie J. Lounsbury, editors. Rescue, Rehabilitation and Release of Marine Mammals: An Analysis of Current Views and Practices: Proceedings of a workshop held in Des Plaines, Illinois, 3–5 December 1991, volume NMFS-OPR 8 of NOAA technical memorandum. NOAA, Washington, DC, 1996. ISBN ???? 65 pp. LCCN ????

Floyd:1996:FPE

[548] R. B. Floyd and A. W. Sheppard, editors. Frontiers of Population Ecology. CSIRO Press, Melbourne, Australia, 1996.

Rhodes:1996:PDE

[549] Olin E. Rhodes, Jr., Ronald K. Chesser, and Michael H. Smith, editors. Population Dynamics in Ecological Space and Time. University of Chicago Press, Chicago, IL, USA, 1996. ISBN 0-226-71057-2 (cloth), 0-226-71058-0 (paper). viii + 388 pp. LCCN QH352 .P628 1996.

Wilkins:1996:N

[550] Robert H. Wilkins and Setti S. Rengachary, editors. Neurosurgery. Mc-Graw-Hill, New York, NY, USA, 1996. ISBN 0-07-079991-1 (set: hard cover), 0-07-070313-2 (vol. 1: hard cover) 0-07-070314-0 (vol. 2: hard cover) 0-07-070315-9 (vol. 3: hard cover). xlv + 4271 + 96 pp. LCCN WL368 N52 1996.

Kunin:1997:BR

[551] W. Kunin and K. Gaston, editors. The Biology of Rarity. Chapman and Hall, Ltd., London, UK, 1997. ISBN ???? ???? pp. LCCN ????

Adler:1997:PIE

[552] Frederick R. Adler. Part I: Ecology and evolution. In Othmer et al. [573], pages 1–2. ISBN 0-13-574039-8. LCCN QH541.15.M3C37

1997.  US$41.00.  URL http://

www.math.utah.edu/books/csmmepcb/
index.html;   http://www.math.utah.
edu/pub/tex/bib/csmmepcb.html;
http://www.phptr.com/.

Ellner:1997:YBY

[553] Stephen P. Ellner. You bet your life: Life-history strategies in fluctuating environments. In Othmer et al. [573], chapter 1, pages 3–24. ISBN 0-13-574039-8. LCCN QH541.15.M3C37

1997.  US$41.00.  URL http://

www.math.utah.edu/books/csmmepcb/
index.html;   http://www.math.utah.
edu/pub/tex/bib/csmmepcb.html;
http://www.phptr.com/.

Holt:1997:ESN

[554] Robert D. Holt and Richard Gomulkiewicz. The evolution of species’ niches: A population dynamic perspective. In Othmer et al. [573], chapter 2, pages 25–50. ISBN 0-13-574039-8. LCCN QH541.15.M3C37

1997.  US$41.00.  URL http://

www.math.utah.edu/books/csmmepcb/

index.html;   http://www.math.utah.
edu/pub/tex/bib/csmmepcb.html;
http://www.phptr.com/.

Diekmann:1997:RME

[555] Odo Diekmann. Reflections on models of epidemics triggered by the case of Phocine Distemper Virus among seals. In Othmer et al. [573], chapter 3, pages 51–59. ISBN 0-13-574039-8. LCCN QH541.15.M3C37

1997.  US$41.00.  URL http://

www.math.utah.edu/books/csmmepcb/
index.html;   http://www.math.utah.
edu/pub/tex/bib/csmmepcb.html;
http://www.phptr.com/.

Nisbet:1997:SRB

[556] R. M. Nisbet, E. McCauley, W. S. C.

Gurney, W. W. Murdoch, and A. M. de Roos. Simple representations of biomass dynamics in structured populations. In Othmer et al. [573], chapter 4, pages 61–79. ISBN 0-13-574039-8. LCCN QH541.15.M3C37

1997.  US$41.00.  URL http://

www.math.utah.edu/books/csmmepcb/
index.html;   http://www.math.utah.
edu/pub/tex/bib/csmmepcb.html;
http://www.phptr.com/.

Tavare:1997:AID

[557] Simon Tavaré. Ancestral inference from DNA sequence data. In Othmer et al. [573], chapter 5, pages 81–96. ISBN 0-13-574039-8. LCCN QH541.15.M3C37

1997.  US$41.00.  URL http://

www.math.utah.edu/books/csmmepcb/
index.html;   http://www.math.utah.
edu/pub/tex/bib/csmmepcb.html;
http://www.phptr.com/.

Lewis:1997:PIC

[558] Mark A. Lewis. Part II: Cell biology. In Othmer et al. [573], pages 97–98. ISBN 0-13-574039-8. LCCN QH541.15.M3C37

1997.  US$41.00.  URL http://

                                        www.math.utah.edu/books/csmmepcb/
                                        index.html;   http://www.math.utah.
                                        edu/pub/tex/bib/csmmepcb.html;
                                        http://www.phptr.com/.

Othmer:1997:STS

[559] Hans G. Othmer. Signal transduction and second messenger systems. In Othmer et al. [573], chapter 6, pages 99–126. ISBN 0-13-574039-8. LCCN QH541.15.M3C37

1997.  US$41.00.  URL http://

                                        www.math.utah.edu/books/csmmepcb/
                                        index.html;   http://www.math.utah.
                                        edu/pub/tex/bib/csmmepcb.html;
                                        http://www.phptr.com/.

Tyson:1997:ECC

[560] John J. Tyson, Kathy Chen, and Bela Novak. The eukaryotic cell cycle: Molecules, mechanisms, and mathematical models. In Othmer et al. [573], chapter 7, pages 127–147. ISBN 0-13-574039-8. LCCN QH541.15.M3C37

1997.  US$41.00.  URL http://

                                        www.math.utah.edu/books/csmmepcb/
                                        index.html;   http://www.math.utah.
                                        edu/pub/tex/bib/csmmepcb.html;
                                        http://www.phptr.com/.

Mackey:1997:MMH

[561] Michael C. Mackey. Mathematical models of hematopoietic cell replication and control. In Othmer et al. [573], chapter 8, pages 149–178. ISBN 0-13-574039-8. LCCN QH541.15.M3C37

1997.  US$41.00.  URL http://

www.math.utah.edu/books/csmmepcb/
index.html;   http://www.math.utah.
edu/pub/tex/bib/csmmepcb.html;
http://www.phptr.com/.

Milton:1997:OMD

[562] John Milton and Jennifer Foss. Oscillations and multistability in delayed feedback control. In Othmer et al. [573], chapter 9, pages 179–198. ISBN 0-13-574039-8. LCCN QH541.15.M3C37

1997.  US$41.00.  URL http://

www.math.utah.edu/books/csmmepcb/
index.html;   http://www.math.utah.
edu/pub/tex/bib/csmmepcb.html;
http://www.phptr.com/.

Sherman:1997:CMP

[563] Arthur Sherman. Calcium and membrane potential oscillations in pancreatic β-cells. In Othmer et al. [573], chapter 10, pages 199–217. ISBN 0-13-574039-8. LCCN QH541.15.M3C37

1997.  US$41.00.  URL http://

www.math.utah.edu/books/csmmepcb/
index.html;   http://www.math.utah.
edu/pub/tex/bib/csmmepcb.html;
http://www.phptr.com/.

Othmer:1997:PIP

[564] Hans G. Othmer. Part III: Physiology. In Othmer et al. [573], pages 219–220. ISBN 0-13-574039-8. LCCN QH541.15.M3C37 1997. US$41.00. URL

http://www.math.utah.edu/books/
csmmepcb/index.html;   http://www.
math.utah.edu/pub/tex/bib/csmmepcb.
html;   http://www.phptr.com/.

Pate:1997:MMM

[565] Edward Pate. Mathematical modeling of muscle crossbridge mechanics. In Othmer et al. [573], chapter 11, pages 221–254. ISBN 0-13-574039-8. LCCN QH541.15.M3C37

1997.  US$41.00.  URL http://

www.math.utah.edu/books/csmmepcb/
index.html;   http://www.math.utah.
edu/pub/tex/bib/csmmepcb.html;
http://www.phptr.com/.

Glass:1997:TPR

[566] Leon Glass. The topology of phase resetting and the entrainment of limit cycles. In Othmer et al. [573], chapter 12, pages 255–276. ISBN 0-13-574039-8. LCCN QH541.15.M3C37

1997.  US$41.00.  URL http://

www.math.utah.edu/books/csmmepcb/
index.html;   http://www.math.utah.
edu/pub/tex/bib/csmmepcb.html;
http://www.phptr.com/.

Krassowska:1997:MIC

[567] Wanda Krassowska. Modeling the interaction of cardiac muscle with strong electric fields. In Othmer et al. [573], chapter 13, pages 277–308. ISBN 0-13-574039-8. LCCN QH541.15.M3C37

1997.  US$41.00.  URL http://

www.math.utah.edu/books/csmmepcb/
index.html;   http://www.math.utah.
edu/pub/tex/bib/csmmepcb.html;
http://www.phptr.com/.

Peskin:1997:FDH

[568] Charles S. Peskin and David M. Mc-Queen. Fluid dynamics of the heart and its valves. In Othmer et al. [573], chapter 14, pages 309–337. ISBN 0-13-574039-8. LCCN QH541.15.M3C37

1997.  US$41.00.  URL http://

www.math.utah.edu/books/csmmepcb/
index.html;   http://www.math.utah.
edu/pub/tex/bib/csmmepcb.html;
http://www.phptr.com/.

Hill:1997:B

[569] N. A. Hill. Bioconvection. In Othmer et al. [573], chapter 15, pages 339–351. ISBN 0-13-574039-8. LCCN QH541.15.M3C37 1997. US$41.00. URL

http://www.math.utah.edu/books/
csmmepcb/index.html;   http://www.
math.utah.edu/pub/tex/bib/csmmepcb.
html;   http://www.phptr.com/.

Adler:1997:AM

[570] Frederick R. Adler. Age-structured models. In Othmer et al. [573], pages 353–356. ISBN 0-13-574039-8. LCCN QH541.15.M3C37 1997. US$41.00. URL

http://www.math.utah.edu/books/
csmmepcb/index.html;   http://www.
math.utah.edu/pub/tex/bib/csmmepcb.
html;   http://www.phptr.com/. Ap-

pendix A.

Lewis:1997:QTO

[571] Mark A. Lewis and Hans G. Othmer.

Qualitative theory of ordinary differential equations. In Othmer et al. [573], pages 357–380. ISBN 0-13-574039-8. LCCN QH541.15.M3C37

1997.  US$41.00.  URL http://

www.math.utah.edu/books/csmmepcb/
index.html;   http://www.math.utah.
edu/pub/tex/bib/csmmepcb.html;
http://www.phptr.com/. Appendix B.

Othmer:1997:IPD

[572] Hans G. Othmer. An introduction to partial differential equations. In Othmer et al. [573], pages 381–386. ISBN 0-13-574039-8. LCCN QH541.15.M3C37

1997.  US$41.00.  URL http://

www.math.utah.edu/books/csmmepcb/
index.html;   http://www.math.utah.
edu/pub/tex/bib/csmmepcb.html;
http://www.phptr.com/. Appendix C.

Othmer:1997:CSM

[573] H. G. Othmer, F. R. Adler, M. A. Lewis, and J. C. Dallon, editors. Case Studies in Mathematical Modeling—Ecology, Physiology, and Cell Biology. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-574039-8. viii + 411 pp. LCCN QH541.15.M3C37

1997.  US$41.00.  URL http://

                                        www.math.utah.edu/books/csmmepcb/
                                        index.html;   http://www.math.utah.
                                        edu/pub/tex/bib/csmmepcb.html;
                                        http://www.phptr.com/.